### v1

# https://www.vrvis.at

Abstract
In this work we address flexibility in deep learning by means of transductive reasoning. For adaptation to new tasks or new data, existing methods typically involve tuning of learnable parameters or even complete re-training from scratch, rendering such approaches unflexible in practice. We argue that the notion of separating computation from memory by the means of transduction can act as a stepping stone for solving these issues. We therefore propose PARMESAN (parameter-free memory search and transduction), a scalable transduction method which leverages a memory module for solving dense prediction tasks. At inference, hidden representations in memory are being searched to find corresponding examples. In contrast to other methods, PARMESAN learns without the requirement for any continuous training or fine-tuning of learnable parameters simply by modifying the memory content. Our method is compatible with commonly used neural architectures and canonically transfers to 1D, 2D, and 3D grid-based data. We demonstrate the capabilities of our approach at complex tasks such as continual and few-shot learning. PARMESAN learns up to 370 times faster than common baselines while being on par in terms of predictive performance, knowledge retention, and data-efficiency.
Keywords:
Transduction Memory Correspondence Matching Fast Learning Continual Learning Few-Shot Learning
1 Introduction
Figure 1: Overview of our parameter-free transduction method for dense prediction tasks. First, a feature extractor $F$ is applied to a query input $x_{q}$ to obtain hidden representations $(h)_{q}^{l}$ for each level $l$. A memory $M$ stores labeled samples along with their hidden representations $(h)_{j}^{l}$. Our approach then performs a hierarchical search in $M$. The search objective is to find both globally and locally similar nodes compared to the query. In each level $l$, we keep the top-$k$ most similar nodes and retrieve their respective child nodes in level $l-1$. Labels $z_{k}$ from most similar memory nodes are retrieved to obtain an initial, raw prediction. Finally, message passing (MP) is applied to obtain a final prediction $y_{q}$. (Simpson images credit: [107, 90, 91, 108])
For training supervised, task-specific deep learning (DL) models, both data and task are usually well-defined in advance. While this setup holds true for many practical problems, a major issue emerges when already existing models require the ability to adapt to both new data and new tasks. Satisfying such a flexibility requirement is non-trivial in practice, adding yet another level of difficulty to a given problem. Although research in fields like continual learning (CL) and few-shot learning (FSL) resulted in many promising approaches, they still suffer from certain restrictions [8, 102] or focus on a specific problem-niche [66]. Moreover, many proposed solutions are often overly complex and require domain experts for deployment and maintenance.
For example, a well established approach is replay-based training in CL, which uses a memory module for knowledge retention [82, 71, 51, 76]. However, replay training turns out to be quite tricky to do in practice due to several reasons. First, models require a high plasticity in order to learn and unlearn continually, which in turn limits their ability to retain knowledge. Forcing the model to be robust, i.e., retain knowledge via replay training, is thus somewhat contradicting the plasticity requirement to learn new concepts. Second, all relevant knowledge has to be incorporated into the model via training, thereby rendering the available memory capacity irrelevant for inference. Third, the balancing of memory samples and new samples for continual training is non-trivial in practice and often leads to the problem of memory overfitting [51, 6, 99]. Finally, replay does not allow to unlearn specific examples easily, e.g., in case they are not needed anymore. Other CL approaches require repeated training from scratch [66], often rendering them infeasible for practical applications.
We argue that simplicity and flexibility are not necessarily mutually exclusive properties of modern DL approaches. In contrast to common, induction-based approaches, where predictions are inferred from general, learned principles, we embrace transduction [21], which is characterized by reasoning from specific training cases to specific test cases.
Our Work: We propose PARMESAN (parameter-free memory search and transduction), an approach that separates computation from memory and extends transduction to complex, high-dimensional labels for dense prediction tasks while at the same time retaining high flexibility. Our method combines several concepts to perform hierarchical correspondence matching with a memory that contains task-specific, labeled data. Rather than requiring to incorporate all knowledge into learnable parameters, the memory in our method extends the capacity of existing models. We further propose a message passing approach to take advantage of correlations within a query sample.
In general, PARMESAN is parameter-free in the sense that it does not have learnable parameters. Instead, learning can be easily and intuitively performed by non-ML experts via memory consolidation, i.e., adding, removing, or modifying memory samples. Our method allows for fast learning and unlearning of individual examples or parts thereof. Unlike many DL approaches for CL and FSL, our method does not require any continual training or fine-tuning. It therefore neither suffers from parameter-induced forgetting and memory overfitting, nor does it require complicated and energy-hungry training by design.
While we apply our method to 2D images in this work, it canonically transfers to other data dimensionalities such as 3D volumes and 1D sequences, as well as being compatible with spatio-temporal data. We do not impose major restrictions on memory size and allow for flexible memory management. Our method can be applied in combination with common DL architectures and is task-agnostic, i.e., transfers to different dense prediction tasks as well as multi-tasking. Our method can be easily integrated into existing DL architectures, thereby enhancing both flexibility and learning speed.
Our Contributions:
  • We propose a flexible, parameter-free transduction method for dense prediction tasks using hierarchical correspondence matching with a memory module.
  • We further propose a parameter-free message passing approach to take advantage of intra-query correlations.
  • We demonstrate that PARMESAN can be used in combination with common feature extractors.
  • We show that our method can be successfully applied to CL and FSL, being on par with commonly used baselines in terms of predictive performance. At the same time, our method learns up to 370 times faster, has stable knowledge retention properties, and is data-efficient. We demonstrate the capabilities of our approach for semantic segmentation and monocular depth estimation on the Cityscapes and JSRT Chest X-ray datasets.
2 Related work
Our method combines concepts from the following fields:
Memory Networks: Recently, the field of memory networks [31, 106, 25, 26, 68] experiences renewed interest. When separating computation and memory [26], some parts of an ML model are responsible for computing, others for storing knowledge. With this notion it is possible to add new information easily, while at the same time being able to treat memory contents as variables. Memory consolidation [23, 3] concerns the strategy for deciding which knowledge to keep in the presence of limited memory resources. Sophisticated memory consolidation methods may include, e.g., novelty search [44], where more informative examples replace less informative examples in memory.
Transduction: Recently, transduction methods experience a revival due to their effectiveness on various classification tasks [2, 37]. One of the most prominent transduction methods is based on k-nearest neighbors (k-NN) [16], which aims to find similar examples w.r.t. common features. Transduction via k-NN was demonstrated to work well for image classification [59], where labeled training samples and their hidden representations are stored in memory. Given a query sample, labels of similar samples are retrieved from memory, followed by a majority voting for prediction. Moreover, k-NN in combination with vision transformers [98, 11] yielded promising results for dense out-of-distribution (OOD) detection [20]. However, making k-NN scalable to dense prediction tasks at high resolutions remains underexplored.
Correspondence Matching: Dense correspondence matching [75, 105, 48] [53, 36, 32, 86, 13] is a fundamental problem in computer vision and often acts as a starting point for various downstream tasks such as optical flow, depth-estimation, etc. In this context, promising methods for improving feature correlation have been proposed [94, 34]. Also, methods for exploiting pyramid features have been introduced for object detection [24, 47] and semantic feature matching [95, 111]. Recently, large-scale foundation models [4] have been applied for one-shot semantic segmentation [49]. Many existing matching approaches use a single reference sample with known correspondences, but typically do not scale well to more than one reference sample. Moreover, most methods rely on specific assumptions commonly found in setups with videos or stereo or multi-view images. In contrast, our proposed method allows matching of arbitrary samples and does not require strong assumptions regarding the presence of correspondences.
Message Passing (MP): Graph neural networks [81, 9, 40] introduced the concept of message passing to DL, where graph nodes are updated depending on received messages from connected nodes. MP was successfully applied, e.g., to various problems in bioinformatics, material science, and chemistry [110, 72].
Continual Learning (CL): The field of CL [93, 73, 92, 61, 8, 102] studies the ability of an ML model to acquire, update, and accumulate knowledge in an incremental manner. This enables models, e.g., to adapt to data-distribution shifts and entirely new tasks. CL is also relevant when frequent re-training is infeasible. Three CL scenarios of increasing difficulty have been proposed [97]: task-incremental (TI), domain-incremental (DI), and class-incremental (CI) CL. The definition of the CI scenario is not limited to classification tasks, but is characterized by the ability of a model to infer which task it is presented with. Prominent approaches for CL include regularization-based methods [41, 46], parameter-isolation methods [54, 80], and replay methods [82, 71, 51, 76]. In contrast to biological neural networks, replay in ML is used for regularizing ML models, where the idea is to learn from new data in tandem with data that is stored in a memory or generated during training. Replay methods can even lead to optimal CL in case of a perfect memory [42].
The notion of knowledge retention and mitigation of forgetting [27] plays a major role in learning. Forgetting is affected by many aspects such as capacity, plasticity, and the quality as well as diversity of training data. The catastrophic forgetting problem in CL [56, 18, 55, 19, 41] refers to learnable parameters adapting to the most recently seen training examples, thereby discarding knowledge from old training examples.
Few-Shot Learning (FSL): Many DL approaches lack the ability to learn efficiently when data is limited. This led to advances in the field of FSL [33], i.e., learning from only a few labeled examples. Available examples are also referred to as support samples. Some methods approach FSL by the means of meta learning [12, 69, 22], where an outer optimization loop learns a strategy for an inner loop, which further adjusts a model to a specific task. The outer loop is performed using large datasets, whereas the inner loop has access to support samples. Another approach to FSL is transfer learning via fine-tuning [15, 60, 100, 1, 45]. Starting from a pre-trained model, learnable parameters or parts thereof are fine-tuned to a specific task using available support samples. Initializing models via pre-training on similar data and tasks typically benefits finding models with high performance after fine-tuning.
Privacy: Common application domains for CL and FSL are autonomous driving and medicine, where data distribution shifts [67] over time are common [87, 64] and data is often scarce. However, storing any kind of personally identifiable data such as patient information in a memory can raise privacy concerns. Even storing hidden representations [29] or detailed, sample-specific labels, raises data-privacy issues due to the possibility of reconstruction attacks [17, 85]. The field of data-free CL [46] has emerged as a viable extension when privacy-preservation is required.
3 Method
We now describe our parameter-free transduction method in more detail (also see Fig. 1). First, we provide an overview of our method and describe the individual components of our model, which consists of a frozen feature extractor, a memory, and a transduction module (Sec. 3.1). Given extracted features of a query input, we perform transductive reasoning by searching for similar features within memory samples and retrieving stored labels (Sec. 3.2). Finally, we take advantage of intra-query correlations via message passing (Sec. 3.3) and further improve predictions by applying test-time augmentation (Sec. 3.4).
3.1 General Setup of PARMESAN
Feature Extraction: Our method requires a pre-trained feature extractor $F$. Let $F$ be a neural network with learnable parameters $\theta$ and consisting of an encoder-like or encoder-decoder-like architecture (e.g., U-Net, ConvNeXt-v2, SwinUNETR, etc.) that delivers a feature pyramid $(h)^{l}$ of hidden representations with levels $l=1,...,n$. We refer to $h^{1}$ as the hidden representation of the first level and $h^{n}$ as the hidden representation of the final level (bottleneck). Individual “pixels” in $(h)^{l}$ with their respective features are referred to as nodes. The total number of nodes for level $l$ is $p^{l}=\prod_{dim}res_{dim}^{l}$, where $dim$ refers to the data dimension and $res^{l}$ to the grid resolutions. Our proposed method is designed with the ability to perform tasks like CL and FSL with frozen parameters $\theta$. One can either pre-train $F$ in a supervised or unsupervised manner, or can use existing pre-trained models. We recommend pre-training $F$ on a dense prediction task in order to enrich features with dense information.
Parameter-free Transduction: Let $X$, $H$, and $Z$ be the input-, latent-, and output spaces, respectively. Densely labeled data is referred to as $(x,z)$ with inputs $x\in X$ and labels $z\in Z$. We define a solver $S:X\rightarrow Z$ as the combination of a feature extractor $F:X\rightarrow H$, a memory $M$, and a parameter-free transduction module $G:H\rightarrow Z$. Given a query input $x_{q}$ and a memory $M$ that stores $m$ labeled samples, $G$ then predicts $y_{q}=G(F(x_{q};\theta),M)$. In essence, $G$ performs hierarchical correspondence matching between a query feature pyramid $(h)_{q}^{l}$ and the memory content, followed by label-retrieval and reduction. We use $M$ for both learning and long-term knowledge retention. Knowledge in $M$ can be modified explicitly, allowing for nuanced and sensible management of what to learn, what to retain, and what to forget. Note that this setup shifts the plasticity-robustness trade-off from learnable parameters $\theta$ with parameter-induced forgetting to $M$ with memory-induced forgetting.
Assumptions: We assume grid-based data structures such as images, volumes, videos, and time series. Although we focus on orthogonal grids in this work, our method can be transferred to irregular, adaptive, and isometric grids. We do not impose any restrictions on label-space growth, such as requiring to pre-define the number of output channels, the number of dimensions, or the numerical data type. When taken to extremes, individual tokens of labels can even represent entire data structures such as sets, series, or graphs.
Privacy Preservation: Our method also extends to privacy-preserving settings, where privacy leaks can be mitigated by storing generated [82] or distilled [103, 28] data in $M$. We do not impose strict requirements for having knowledge about labels, e.g., class segments, and their distribution, e.g., class abundances. This setup even allows for keeping memory labels private if needed.
3.2 Hierarchical Correspondence Matching
We initialize our model by filling $M$ with $m$ labeled samples consisting of $(x,z)_{j}$ as well as their respective feature pyramids $(h)_{j}^{l}$, where $j=1,...,m$ and $l=1,...,n$. Leaf nodes in level $l=0$ refer to memory labels $z_{j}$. Next, we define connectivity kernels which define children-parent relations across different levels. Kernels remain fixed for every sample and for every level. For more details, see Appendix 0.A.
Some neural architectures may not provide desired feature pyramids in the sense that resolutions are halved at every level or that bottlenecks have small resolutions. In these cases, connectivity kernels can either be adjusted accordingly, or missing representations can be inserted by interpolation or extrapolation of neighboring representations.
Algorithm 1 PARMESAN model inference
$dim$, $\phi$triangleright hyperparameters
$x_{q},(h)_{j}^{l},z_{j}\ ;j=1,...,m\ ;l=1,...,n$triangleright input
$(h)_{q}^{l}\leftarrow F(x_{q};\theta)\ ;l=1,...,n$triangleright encode query
init (in parallel for all $p^{n}$ query nodes):
$k\leftarrow m$
$id_{k,par}\leftarrow p^{n}\times[0,1,2,...,k]$triangleright parent indices
$s_{k,par}\leftarrow p^{n}\times[1.,1.,1.,...,1.]$triangleright parent similarities
memory search (in parallel for all $p^{l}$ query nodes):
for $l=n,1$ do
     $id_{comp}\leftarrow get\_children(l+1,id_{k,par})$triangleright [$p^{l}\times k\cdot 2^{dim}$]
     $s^{l}\leftarrow calc\_sim(h_{q}^{l},h_{j}^{l}[id_{comp}])$triangleright [$p^{l}\times k\cdot 2^{dim}$]
     $s_{acc}^{l}\leftarrow s^{l}\cdot s_{k,par}$triangleright update, [$p^{l}\times k\cdot 2^{dim}$]
     if $l>1$ then
         $k\leftarrow\lfloor\phi\cdot k\rfloor$triangleright reduce
     end if
     $id_{keep}\leftarrow get\_topk(s_{acc}^{l},k)$triangleright find top-$k$, [$p^{l}\times k$]
     $id_{k,par}\leftarrow id_{comp}[id_{keep}]$triangleright update, [$p^{l-1}\times k$]
     $s_{k,par}\leftarrow s_{acc}^{l}[id_{keep}]$triangleright update, [$p^{l-1}\times k$]
end for
label-retrieval and reduction:
$id_{z}\leftarrow get\_children(l=1,id_{k,par})$
$y_{q,raw}\leftarrow sum(softmax(s_{k,par})_{k}\cdot z_{j}[id_{z}])_{k}$
return $y_{q,raw}$triangleright raw prediction
We formulate the transductive forward pass of our method in Algorithm 1. First, a query $x_{q}$ is encoded via $F$ to obtain $(h)_{q}^{l}$. Then, $G$ takes $(h)_{q}^{l}$ together with memory feature pyramids $(h)_{j}^{l}$ and their respective labels $z_{j}$ as inputs. The overall objective is searching for both globally and locally similar nodes in $M$. Starting from $l=n$, the search objective in each level is first comparing queries $h_{q}^{l}$ with keys $h_{j}^{l}$, followed by retrieving the children of the top-$k$ most similar memory nodes. We use the cosine similarity between nodes’ hidden features as our similarity measure. The accumulated similarity $s_{acc}^{l}$ between two nodes at level $l$ is defined as the product of their similarities $s^{l}$ at level $l$ and the similarities $s_{k,par}$ of their respective parents across all higher levels:
$s_{acc}^{l}=\prod_{u=l}^{n}sim(h_{q}^{u},h_{j}^{u})=\prod_{u=l}^{n}\frac{h_{q}^{u}\cdot h_{j}^{u}}{\lVert h_{q}^{u}\rVert\ \lVert h_{j}^{u}\rVert}$ (1)
For each of the $p^{l}$ query nodes, we use $s_{acc}^{l}$ to identify the top-$k$ most similar nodes in level $l$ and retrieve their respective children in level $l-1$. This implies that the top-$k$ most similar nodes in level $l$ become parent nodes for level $l-1$. Children nodes in level $l-1$ are required for processing level $l-1$, thereby forming a hierarchy across levels. The hyperparameter $\phi\in(0.0,1.0]$ controls the number $k$ of top-$k$ nodes to retain for every query node in each level during search. Effectively, $\phi$ is the rate of reduction for $k$ in each level. A variable number of nodes can be retained in different memory samples.
Considering the data dimension $dim$ and the grid resolutions $res^{l}$, exhaustive search would result in a computational complexity of $\mathcal{O}((\prod_{dim}res_{dim}^{l})^{2})$ for level $l$. In contrast, our approach performs search locally and has a computational complexity of only $\mathcal{O}(n_{ch}\cdot(\prod_{dim}res_{dim}^{l}))$ with $n_{ch}$ being the number of child nodes in level $l+1$. Also, since search is performed independently for every query node, our approach can be highly parallelized in practice and is thus scalable to high resolutions and high-dimensional data. Moreover, since $h^{l}$ are strongly correlated across levels, there is a high chance that similarities to query nodes are also strongly correlated across levels. In other words, searching locally pre-filters nodes that are unlikely to be among the top-$k$ most similar nodes as search progresses to the next level.
Memory search results in a dense correspondence map between a query $x_{q}$ and, for every query node, their top-$k$ most similar memory nodes $i_{k,par}$. We use these matches to retrieve task-specific labels $z_{k}$ from $M$. To obtain a raw prediction $y_{q,raw}$, we use the softmax-normalized similarities $s_{k,par}$ of all matches to obtain a weighted average of retrieved labels.
3.3 Intra-Query Transduction via Message Passing
Up until this point we did not exploit any intra-query correlations such as applying geometric priors or enforcing smoothness of predictions. Strong local correlations are common in grid-based data and should be exploited if possible. The simplest variant of intra-query transduction is Gaussian blurring, where local smoothness over neighboring nodes is enforced. However, blurring is not content-aware and can harm performance at certain regions, e.g., at sharp corners or small objects. We therefore propose a content-aware, parameter-free message passing (MP) approach to account for local correlations within the query and achieve pixel-exact predictions. We re-use our hierarchical search without memory content to find the top-$\kappa$ nearest neighbor nodes within the query sample itself. Similarity scores between a query node $i$ and its nearest neighbors $j$ are referred to as edges $e_{ij}$. We then perform MP to refine raw predictions $y_{q,raw}$. Considering a query node state $y_{q,i}$, we define the equations for the edges $e_{ij}$, the aggregated message $\hat{y}_{q,i}$, and the node update from step $t\rightarrow t+1$ as
$e_{ij}=\frac{exp(s_{acc,j}^{1})}{\sum_{j=1}^{\kappa}exp(s_{acc,j}^{1})}\ ,$ (2)
$\hat{y}_{q,i}=\sum_{j=1}^{\kappa}e_{ij}\ y_{q,j}^{t}\ ,$ (3)
$y_{q,i}^{t+1}=(1-\lambda)\ y_{q,i}^{t}+\lambda\ \hat{y}_{q,i}\ .$ (4)
The hyperparameter $\lambda\in(0.0,1.0]$ controls the overall strength of MP per step. We apply Eq. 3 and Eq. 4 repeatedly for all query nodes in level $l=0$ until convergence [113], i.e., when $||y_{q,i}^{t+1}-{y}_{q,i}^{t}||\rightarrow 0$. The resulting prediction $y_{q}$ then contains less artifacts and is spatially more consistent compared to $y_{q,raw}$.
3.4 Test-Time Augmentation
Similar to recent work [36], we perform test-time augmentation (TTA) to resolve small errors. We apply multiple forward passes with down-scaled and reflection-padded variants of $x_{q}$. Registration of down-scaled variants with the original variant is achieved by up-scaling and center-cropping the respective predictions. We obtain a final prediction $y_{q}$ by averaging over all passes.
3.5 Fast Learning
PARMESAN is designed to learn fast in the sense that it does not require training of learnable parameters $\theta$ to perform complex tasks such as CL or FSL. In contrast to other approaches, our method only requires memory consolidation, i.e., saving examples in $M$ and extracting their respective feature pyramids $(h)_{j}^{l}$ in a single forward pass using $F$ with pre-trained, frozen $\theta$. For a given hardware setup, learning speed can be split up into required wall-clock times $\tau_{p}$ and $\tau_{l}$ for pre-training and learning, respectively. We define $\tau_{p}$ on a per-dataset basis and $\tau_{l}$ on a per-sample basis. The total required wall-clock time to learn $n$ training samples is then $T=\tau_{p}+n\cdot\tau_{l}$.
4 Experiments and Results
We demonstrate the capabilities of our method in three groups of experiments:
A0 – A18:
ablation studies for semantic segmentation and monocular depth estimation to demonstrate flexibility (Sec. 4.1, Tab. 1, and Tab. 2),
B0 – B7:
CL experiments for studying knowledge retention properties of our method (Sec. 4.2, Tab. 3, and Tab. 4),
C0 – C7:
FSL experiments to investigate data-efficiency (Sec. 4.3 and Tab. 5).
More details regarding applied models, training and evaluation, as well as qualitative results for all experiments are provided in Sec. 0.B.2.
Our main goal is to improve flexibility and adaptation capabilities of DL models while simplifying and speeding up learning for CL and FSL. Problem-specific methods are expected to outperform more general methods such as ours [57, 104, 101].
Datasets: We use the Cityscapes (CITY) benchmark dataset [7] for semantic segmentation and monocular depth estimation. Semantic segmentation masks comprise fine and coarse labels, whereas disparity maps are provided for depth estimation. We also use the JSRT Chest X-ray dataset [84] for semantic segmentation. We refer to Sec. 0.B.1 for further details.
Method: We employ a U-Net [77] variant as our default DL model and feature extractor. For experiments A11 and A14 we use ConvNeXt Tiny [50] with $\theta$ trained on ADE20K [112]. For all experiments that require $M$, we select memory samples randomly from the training split while keeping the random generation process fixed to ensure comparability. For PARMESAN, we set $m~{}=~{}256$, $\phi=0.5$, apply message passing (MP) and test-time augmentation (TTA), and use both pre-trained U-Net encoder and decoder for feature extraction unless stated otherwise. We use Gaussian blurring with $\sigma=5$ pixels in experiment A2. CL and FSL are performed by memory consolidation of training samples. To obtain memory samples for the privacy-preserving experiment A6, we additionally apply data distillation via linear mixup of two random training samples [109, 43, 83, 70].
Baselines: Joint training (JOINT) refers to training on all available training data, thus representing the default supervised learning setup and the upper bound for CL [97]. Fine-tuning (FT) is a simple, yet strong baseline for FSL [10]. Due to catastrophic forgetting, however, FT is regarded as the lower bound for CL [97]. We further employ memory-based approaches such as classical replay (REPLAY) [82, 71, 51, 76] and greedy-sampling dumb learning (GDUMB) [66] as additional baselines for CL. With GDUMB, training is performed from scratch for each individual CL step, thereby exclusively using memory samples. For consistent and fair comparisons to our method, we implement all baselines and adjust method-specific hyperparameters to our data setup.
Evaluation: We use the mean Intersection over Union of classes ($\text{mIoU}_{\text{cl}}$) and categories ($\text{mIoU}_{\text{cat}}$) as evaluation metrics for semantic segmentation and Root-Mean-Square Error (RMSE) in meters for depth estimation. We report best performances on the CITY validation set and on the JSRT test set, since labels for the CITY test set are not publicly available.
4.1 Ablation Studies
We perform ablation studies of various components to analyze the impact on learning speed and predictive performance, as well as demonstrating the flexibility of our method with different architectures of $F$. We summarize our results in Tab. 1 (segmentation) and Tab. 2 (depth) and visualize predictions in Fig. 2 (CITY, segmentation), Fig. 3 (CITY, depth), and Fig. 4 (JSRT, segmentation).
Learning with our method can be done in less than 0.1 seconds per sample (see column $\tau_{l}$ in Tab. 1 and Tab. 2), thereby mitigating the demand for additional computational costs when attaching PARMESAN to existing models. Our method outperforms JOINT on CITY when trained on coarse labels (A9, A10) and evaluated on fine labels. As expected, our method under-performs JOINT when using fine labels (A0, A1), on JSRT (A12, A13, A14), and for depth estimation (A15, A16). Using only the encoder (A7) or using an entirely untrained, randomly initialized model (A8) yields promising results already, indicating that our transduction module $G$ imposes useful inductive biases. PARMESAN can be successfully used in combination with a ConvNeXt feature extractor on both CITY and JSRT when trained on ADE20K [50] (A11, A14). MP and TTA (A1, A3, A4, A16, A17) considerably improve performance over raw predictions (A5, A18). As expected, content-aware MP (A1) also outperforms non-content-aware Gaussian blurring (A2). Comparing panels A1 and A5 in Fig. 2, MP and TTA succeed in reducing small artifacts while retaining sharp corners of predicted segments. In contrast, blurring typically comes at the cost of missing small segments and sharp corners. Privacy-preservation by data distillation (A6) slightly decreases performance compared to the non-private setup (A1).
Figure 2: Model predictions and analysis of PARMESAN. Left: We study top-$1$ nearest neighbors from experiment A1. The idx panel visualizes indices of retrieved nearest neighbor nodes from $M$. Similar colors indicate that nodes are retrieved from close-by memory locations, most likely from within the same samples. The sim panel visualizes the respective similarity scores, where bright regions refer to high similarities. We also show 3 query pixels and their matches in $M$, indicating both global and local semantic correspondence. Right: PARMESAN can be applied flexibly (A5, A11) and perform complex tasks like CL (B5, B7) and FSL (C3, C7) via memory consolidation.
Figure 3: Model predictions for monocular depth estimation experiments.
Figure 4: Model predictions for semantic segmentation experiments using JSRT data.
In general, we observe global and local semantic correspondence of retrieved nearest neighbors from $M$, allowing for future work on various topics related to interpretability and explainability of predictions. Moreover, since every retrieved node comes with its similarity score w.r.t. the query, our method provides pixel-exact indications for uncertainty. As visible in the sim panel in Fig. 2, similarity scores seem to be correlated with the heterogeneity of regions in the image.
Table 1: Results for ablation studies showing semantic segmentation performance for classes ($\text{mIoU}_{\text{cl}}$) and categories ($\text{mIoU}_{\text{cat}}$). Column “$F$ pre-train/train“ indicates if the method uses a pre-trained feature extractor $F$ and if any additional training of learnable parameters $\theta$ is required. $\tau_{p}$ and $\tau_{l}$ are wall-clock times for pre-training and learning, respectively. Experiment IDs marked with * use ConvNeXt pre-trained on ADE20K [50] for feature extraction.
ID Train data (#samples) Method & Setup $F$ pre-train/train $\tau_{p}$ [h] $\tau_{l}$ [s] $\text{mIoU}_{\text{cl}}$ $\text{mIoU}_{\text{cat}}$
A0 CITY, fine (2975) JOINT no / yes 0 3.72 82.5 87.4
A1 CITY, fine (256) ours A0 / no 3.08 0.09 80.0 85.5
A2 CITY, fine (256) ours, no MP, blur A0 / no 3.08 0.09 79.3 84.8
A3 CITY, fine (256) ours, no TTA A0 / no 3.08 0.09 77.7 83.5
A4 CITY, fine (256) ours, no MP A0 / no 3.08 0.09 77.9 83.2
A5 CITY, fine (256) ours, no MP, no TTA A0 / no 3.08 0.09 75.7 81.6
A6 CITY, fine (256) ours, private A0 / no 3.08 0.10 77.6 83.2
A7 CITY, fine (256) ours A0 (E only) / no 3.08 0.08 72.8 79.1
A8 CITY, fine (256) ours no / no 0 0.09 56.1 63.5
A9 CITY, coarse (19998) JOINT no / yes 0 0.43 49.6 50.6
A10 CITY, coarse (256) ours A9 / no 2.41 0.09 50.7 52.5
A11* CITY, fine (256) ours E only / no - 0.09 74.5 81.4
A12 JSRT (199) JOINT no / yes 0 23.39 93.3 -
A13 JSRT (199) ours, $m\!=\!199$, $\phi\!=\!0.8$ A12 / no 1.29 0.02 91.9 -
A14* JSRT (199) ours, $m\!=\!199$, $\phi\!=\!0.8$ E only / no - 0.02 87.2 -
Table 2: Results for ablation studies showing depth estimation performance.
ID Train data (#samples) Method & Setup $F$ pre-train/train $\tau_{p}$ [h] $\tau_{l}$ [s] RMSE [m]
A15 CITY, disparity (2975) JOINT no / yes 0 6.71 10.5
A16 CITY, disparity (256) ours A15 / no 5.55 0.03 11.9
A17 CITY, disparity (256) ours, no MP A15 / no 5.55 0.03 12.5
A18 CITY, disparity (256) ours, no MP, no TTA A15 / no 5.55 0.03 13.5
4.2 Continual Learning
[…]
Table 3: Results for CL experiments. Performance is measured on all classes after the final CL step (i.e., after 4 steps). Our method is on-par with CL baselines while learning substantially faster.
ID Train data (#samples) Method & Setup $F$ pre-train/train $\tau_{p}$ [h] $\tau_{l}$ [s] $\text{mIoU}_{\text{cl}}$ $\text{mIoU}_{\text{cat}}$
B0 CITY, fine (2975) JOINT (upper bound) no / yes 0 20.62 81.7 86.6
B1 CITY, fine (2975) FT (lower bound) no / yes 0 12.76 4.0 4.3
B2 CITY, fine (2975) REPLAY, $m\!=\!256$ no / yes 0 5.11 56.2 64.2
B3 CITY, fine (256) GDUMB, $m\!=\!256$ no / yes 0 110.9 74.4 80.3
B4 CITY, fine (256) ours B0 ($1^{\text{st}}$ step) / no 3.07 0.32 66.0 70.7
B5 CITY, fine (256) ours B3 ($1^{\text{st}}$ step) / no 1.42 0.32 65.4 71.0
B6 CITY, fine (256) GDUMB, $m\!=\!256$ A9 / yes 2.41 63.13 76.9 82.1
B7 CITY, fine (256) ours A9 / no 2.41 0.32 75.4 81.3
[…]
REPLAY (B2) GDUMB (B3) ours (B4) ours (B5) GDUMB (B6) ours (B7)
[…]
4.3 Few-Shot Learning
For FSL, we pre-train $F$ on coarse labels (A9). We then perform experiments with 10, 8, 6, 4, and 2 fine-labeled support samples. FSL via FT is performed by fine-tuning $\theta$ on support samples. To study FT under realistic conditions, we divide the support sets into 2 splits with the latter containing 4 samples used for early stopping. For FSL with PARMESAN, we fill $M$ with all support samples. A low number of support samples can lead to a high variance in performance. We thus repeat each experiment 5 times with randomly drawn support samples, where we fix the random generation process across experiments to ensure comparability. We report average performances on the full validation set in Tab. 5.
Table 5: Results for FSL experiments, averaged over 5 repetitions. We use a feature extractor pre-trained on coarse labels and a support set containing fine labels. Our method learns fast and is data-efficient especially for small support sets.
ID Train data (#samples) Method & Setup $F$ pre-train/train $\tau_{p}$ [h] $\tau_{l}$ [s] $\text{mIoU}_{\text{cl}}$ $\text{mIoU}_{\text{cat}}$
C0 CITY, fine (10, train on 6) FT A9 / yes 2.41 32.29 71.7 76.4
C1 CITY, fine (8, train on 4) FT A9 / yes 2.41 69.53 71.1 75.9
C2 CITY, fine (6, train on 2) FT A9 / yes 2.41 225.6 69.5 74.7
C3 CITY, fine (10) ours, $m\!=\!10$, $\phi\!=\!1$ A9 / no 2.41 0.17 72.2 78.2
C4 CITY, fine (8) ours, $m\!=\!8$, $\phi\!=\!1$ A9 / no 2.41 0.21 71.1 77.5
C5 CITY, fine (6) ours, $m\!=\!6$, $\phi\!=\!1$ A9 / no 2.41 0.26 69.9 76.2
C6 CITY, fine (4) ours, $m\!=\!4$, $\phi\!=\!1$ A9 / no 2.41 0.34 66.9 73.3
C7 CITY, fine (2) ours, $m\!=\!2$, $\phi\!=\!1$ A9 / no 2.41 0.61 58.3 65.6
FSL with PARMESAN can be done up to 370 times faster w.r.t. FT (C2, C7) and remains on par in terms of predictive performance for equal support set sizes (C0-C5). Our method is data-efficient in the sense that is able to utilize the entire support set at inference and that it can be applied in combination with a small support set (C6, C7).
4.4 Limitations
Depending on the choice of hyperparameters such as $m$ and $\phi$, PARMESAN might require more computational resources and longer inference times compared to learned task-heads. Memory consumption and inference speed can potentially be further improved by introducing sparsity to stored feature pyramids, i.e., reducing the number of unique patterns per memory sample. Moreover, choosing a suitable feature extractor might require taking a trade-off between a general-purpose model versus a data- and task-specific model.
5 Conclusion
In this work we proposed PARMESAN, a flexible, parameter-free approach that separates computation from memory and leverages transduction for solving dense prediction tasks. We also proposed a message passing approach to take advantage of intra-query correlations. Our method can perform complex tasks such as CL and FSL in an easy, intuitive, and optionally private manner without any continuous training of learnable parameters.
We demonstrated that our approach learns substantially faster than common baselines while achieving competitive predictive performance, showing stable knowledge retention, and being data-efficient.
Future Work: Possible future work includes leveraging self-supervised feature extractors, studying OOD detection capabilities, as well as investigating memory-related topics such as dataset distillation. Performance might be further increased by implementing algorithmic or methodical improvements. This includes, e.g., using explorative memory search instead of our current greedy approach, as well as applying novelty search instead of random sampling for memory consolidation. Finally, PARMESAN might be beneficial to other fields such as reinforcement learning, where fast adaptation can be important.
Supplementary Material
Philip Matthias Winter, Maria Wimmer, David Major, Dimitrios Lenis, Astrid Berg, Theresa Neubauer, Gaia Romana De Paolis, Johannes Novotny, Sophia Ulonska, Katja Bühler
Appendix 0.A Method
Initialization: We now describe the initialization of $G$ in more detail. First, we need to initialize hyperparameters such as the number of data dimensions $dim$, the number of memory samples $m$ and the resolutions $(res)^{l}$ for all levels. We then define connectivity kernels between nodes of different levels, which remain fixed for all samples. Each node in level $l$ has a parent node in level $l+1$. For the highest level $l=n$, all nodes have a single root node as their parent. Each node in level $l$ typically has several child nodes in level $l-1$, e.g., $n_{ch}=2^{dim}$. Leaf nodes in level $l=0$ refer to memory labels. We assume that resolutions of $h^{1}$, $x$, and $z$ match, implying that each node in level $l=1$ has exactly one child node in $l=0$. We initialize parent indices $i_{k,par}$ such that hierarchical search has access to the full memory content. We initialize parent similarity scores $s_{k,par}$ with a value of 1.0 for all nodes.
Technical Aspects: In order to avoid increasingly exhaustive search due to a rapidly growing number of nodes, we set $\phi$ such that hardware memory requirements remain quasi-constant across levels. For example, $\phi=0.5$ for 2D data in combination with doubling the channel dimension in each level yields quasi-constant hardware memory usage. The same is true for 3D data and $\phi=0.25$. This is especially relevant in case of high-dimensional data and high resolutions, where available hardware often imposes hard limits on the choice of hyperparameters. Naturally, smaller $m$ allow for larger $\phi$. For depth estimation with PARMESAN, we ignore invalid labels in $M$ by manually setting their respective similarity scores $s_{k,par}=-100.0$ before applying the softmax operation.
We implemented all methods and experiments using the PyTorch framework [63]. We apply the exact same pre- and postprocessing to retain comparability. We conduct all experiments on a single NVIDIA GeForce RTX 3090 (24 GB) graphics card. PARMESAN has inference times in the order of 1 second per image for $m=256$, $\phi=0.5$, and an input resolution of 512$\times$1024 pixels. By default, we store memory content on RAM and transfer representations to GPU when needed for hierarchical search. In case of small GPU RAM, the hierarchical search can be chunked into smaller pieces via windowing and performed sequentially without compromising performance. For MP, all operations can be done on GPU.
Appendix 0.B Experiments
0.B.1 Datasets
From the Cityscapes semantic segmentation benchmark dataset (CITY) [7], we use the 5000 images with fine semantic and disparity annotations, and the 19998 images with coarse semantic annotations (see Fig. 5). We use the official splits for training and validation with 2975 and 500 samples, respectively. Since labels for the 1525 test samples are not publicly available, we report best performances achieved on the validation split. For JSRT [84], a 80:10:10 split is applied, resulting in 199, 24, and 24 samples for training, validation, and testing, respectively. For CITY, we use 34 semantic classes with 8 categories for training and the official 19 classes for evaluation. For JSRT, we use 6 classes for training and evaluation: background, left lung, right lung, heart, left clavicle, and right clavicle. We rescale all samples to a resolution of 512$\times$1024 pixels for CITY and 512$\times$512 pixels for JSRT.
For CITY disparity maps, we observe that some pixels near image borders are labeled incorrectly. We manually mask out these erroneous labels and treat them as invalid measurements.
Figure 5: Annotations of Cityscapes dataset (original image, depth map, fine semantic map, coarse semantic map). While coarse annotations only outline the most prominent objects via polygons, pixel-exact fine annotations are regarded as the ground truth.
0.B.2 Models, Methods, Training and Evaluation
Models: We use a U-Net [77] variant with residual blocks [30] and instance normalization [96] as our default architecture, resulting in 9 levels (excluding the label-level $l=0$) with respective feature map resolutions of [512[512[ 512, $256$, $128$, $64$, $32$, $16$, $8$, $4$, 2]2]2 ] pixels per dimension. The respective number of feature maps are [6[6[ 6, $12$, $24$, $48$, $96$, $192$, $384$, $768$, 1536]1536]1536 ], resulting in a total of 73.15m learnable parameters.
For ablations studies, we additionally use a ConvNeXt Tiny with pre-trained weights for feature extraction [50]. Pre-trained weights were obtained by training ADE20K [112] for semantic segmentation. The model comprises 4 levels with feature map resolutions of $[128,64,32,16]$ and corresponding number of feature maps $[96,192,384,786]$. To be able to use it with our transduction module, we complete the feature pyramid, i.e., add missing resolutions, via interpolation and extrapolation of neighboring representations. We reduce the number of feature maps in high resolutions by a factor of 4 to meet our hardware limitations (see paragraph “Training“ below) and keep the number of feature maps constant for missing lower resolutions. This results in $[6,24,96,192,384,768,768,768,768]$ feature maps at the respective resolutions of $[512,256,128,64,32,16,8,4,2]$.
Method: For memory search, we set $n_{ch}=2^{dim}$ for all levels. For MP, we set $\lambda=1$, $k=16$. Convergence is typically achieved within 16 MP steps. For TTA, we apply three separate forward passes and down-scale inputs by factors $s=\{0.8,0.9,1\}$, where $s=1$ implies no transformation.
[…]
Evaluation: We use exponential moving average models [65, 79, 89] with a rate of 0.99 for evaluation. Early stopping [58] with a patience of 3 and a validation verbosity of 3000 update steps is used (400 for JSRT). Early stopping in CL experiments only depends on the performance on available classes at the respective CL step. We use a validation verbosity of 30 update steps for FSL FT experiments due to small support sets.
We use the Intersection over Union (IoU) [35, 88] as evaluation metric for semantic segmentation, which is defined as the ratio between the size of the intersection and the size of the union of two finite sample sets, measuring their similarity. In this work, we use IoU to measure predictive performance of a model for the task of pixel-wise, multi-class classification, i.e., semantic segmentation. We employ the standard evaluation procedure as defined in [7]. Specifically, we apply the mean IoU ($\text{mIoU}_{\text{}}$), where scores are calculated globally without favoring any class in particular. This variant is also referred to as “micro”-averaging.
Results: We show qualitative prediction results for all ablation studies, CL experiments, and FSL experiments in Fig. 6, Fig. 7, and Fig. 8, respectively.
0.B.3 Continual Learning Setup
[…]
In contrast to image classification, samples can contain classes that belong to different CL steps, which is referred to as overlapping labels. For the CI-CL scenario, this requires introducing a dedicated background class to avoid data-leakage between different CL steps. Therefore, non-available classes are re-labeled as background in each CL step. This implies that the background class is different for different CL steps. As proposed in recent work [5], we ignore background pixels for validation.
Our method supports any combination of the three CL scenarios TI, DI, and CI. Although we focus on studying the offline CL scenario in this paper, our method can also be used for online CL, where one does not have access to the entire dataset of a specific task in order to learn continually.
In our experiments, we apply classical replay (REPLAY) as a baseline method. In order to keep labels consistent for replay training, we employ an adaptive masking approach. For samples of the current CL step $i$, we mask out all background pixels, i.e., pixels which do not belong to currently active classes. In contrast, for memory samples we mask out all pixels which belong to currently active classes, i.e., keeping pixels of old classes and the background. Training minibatches for REPLAY consist of standard samples in addition to replay samples, which we select randomly from $M$. For REPLAY, we observe that using one or more replay samples from $M$ per update step leads to strong memory overfitting, which in turn leads to reduced validation performance. We achieve best results with one replay sample every other minibatch.
Figure 6: Model predictions for all ablation studies with two different query inputs.
Figure 7: Model predictions for all CL experiments with two different query inputs.
Figure 8: Model predictions for all FSL experiments with two different query inputs.

### v2

@@
 Abstract
-In this work we address flexibility in deep learning by means of transductive reasoning.
-For adaptation to new tasks or new data, existing methods typically involve tuning of learnable parameters or even complete re-training from scratch, rendering such approaches unflexible in practice.
+This work addresses flexibility in deep learning by means of transductive reasoning.
+For adaptation to new data and tasks, e.g., in continual learning, existing methods typically involve tuning learnable parameters or complete re-training from scratch, rendering such approaches unflexible in practice.
 We argue that the notion of separating computation from memory by the means of transduction can act as a stepping stone for solving these issues.
-We therefore propose PARMESAN (parameter-free memory search and transduction), a scalable transduction method which leverages a memory module for solving dense prediction tasks.
-At inference, hidden representations in memory are being searched to find corresponding examples.
-In contrast to other methods, PARMESAN learns without the requirement for any continuous training or fine-tuning of learnable parameters simply by modifying the memory content.
-Our method is compatible with commonly used neural architectures and canonically transfers to 1D, 2D, and 3D grid-based data.
-We demonstrate the capabilities of our approach at complex tasks such as continual and few-shot learning.
-PARMESAN learns up to 370 times faster than common baselines while being on par in terms of predictive performance, knowledge retention, and data-efficiency.
+We therefore propose PARMESAN (parameter-free memory search and transduction), a scalable method which leverages a memory module for solving dense prediction tasks.
+At inference, hidden representations in memory are being searched to find corresponding patterns.
+In contrast to other methods that rely on continuous training of learnable parameters, PARMESAN learns via memory consolidation simply by modifying stored contents.
+Our method is compatible with commonly used architectures and canonically transfers to 1D, 2D, and 3D grid-based data.
+The capabilities of our approach are demonstrated at the complex task of continual learning.
+PARMESAN learns by 3-4 orders of magnitude faster than established baselines while being on par in terms of predictive performance, hardware-efficiency, and knowledge retention.
 Keywords:
-Transduction Memory Correspondence Matching Fast Learning Continual Learning Few-Shot Learning
+Transduction Memory Correspondence Matching Fast Learning Continual Learning
 1 Introduction
 Figure 1: Overview of our parameter-free transduction method for dense prediction tasks.
-First, a feature extractor $F$ is applied to a query input $x_{q}$ to obtain hidden representations $(h)_{q}^{l}$ for each level $l$.
-A memory $M$ stores labeled samples along with their hidden representations $(h)_{j}^{l}$.
-Our approach then performs a hierarchical search in $M$.
-The search objective is to find both globally and locally similar nodes compared to the query.
-In each level $l$, we keep the top-$k$ most similar nodes and retrieve their respective child nodes in level $l-1$.
-Labels $z_{k}$ from most similar memory nodes are retrieved to obtain an initial, raw prediction.
-Finally, message passing (MP) is applied to obtain a final prediction $y_{q}$.
-(Simpson images credit: [107, 90, 91, 108])
-For training supervised, task-specific deep learning (DL) models, both data and task are usually well-defined in advance.
-While this setup holds true for many practical problems, a major issue emerges when already existing models require the ability to adapt to both new data and new tasks.
-Satisfying such a flexibility requirement is non-trivial in practice, adding yet another level of difficulty to a given problem.
-Although research in fields like continual learning (CL) and few-shot learning (FSL) resulted in many promising approaches, they still suffer from certain restrictions [8, 102] or focus on a specific problem-niche [66].
-Moreover, many proposed solutions are often overly complex and require domain experts for deployment and maintenance.
-For example, a well established approach is replay-based training in CL, which uses a memory module for knowledge retention [82, 71, 51, 76].
-However, replay training turns out to be quite tricky to do in practice due to several reasons.
-First, models require a high plasticity in order to learn and unlearn continually, which in turn limits their ability to retain knowledge.
-Forcing the model to be robust, i.e., retain knowledge via replay training, is thus somewhat contradicting the plasticity requirement to learn new concepts.
-Second, all relevant knowledge has to be incorporated into the model via training, thereby rendering the available memory capacity irrelevant for inference.
-Third, the balancing of memory samples and new samples for continual training is non-trivial in practice and often leads to the problem of memory overfitting [51, 6, 99].
+A feature extractor $F$ is applied to a query $x_{q}$ to obtain hidden representations $(h)_{q}^{l}$.
+A memory $M$ stores labeled samples along with their representations $(h)_{j}^{l}$.
+Our approach then performs a hierarchical search in $M$, where the objective is to find globally and locally similar nodes w.r.t.
+$x_{q}$.
+In each level $l$, we keep the top-$k$ most similar nodes and retrieve their child nodes in level $l-1$.
+Labels from most similar nodes are retrieved to obtain a raw prediction.
+Then, message passing (MP) is used to get a final prediction $y_{q}$.
+For training task-specific deep learning (DL) models, both data and task are usually well-defined in advance.
+While this setup holds true in most cases, a major issue emerges when already existing models are required to adapt to new data and new tasks.
+Satisfying such a flexibility requirement is non-trivial, increasing the difficulty for solving a given problem.
+Although research in continual learning (CL) resulted in promising approaches, they still suffer from certain restrictions [9, 91] or focus on specific problem-niches [60].
+Moreover, proposed solutions are often unstable [7, 40, 39] and require domain experts for deployment and maintenance.
+For example, a well established CL approach is replay training, which uses a memory module for knowledge retention [75, 63, 46, 68].
+However, replay training turns out to be challenging in practice due to several reasons.
+First, models require high plasticity to learn and unlearn continually, which in turn limits their ability to retain knowledge.
+Forcing models to be robust, i.e., retain knowledge via replay training, is thus somewhat contradicting the plasticity requirement for learning new concepts.
+Second, all relevant knowledge has to be incorporated into model parameters via training, thereby rendering the available memory capacity irrelevant for inference.
+Third, the balancing of memory samples and new samples for training is non-trivial in practice and often leads to memory overfitting [46, 6, 90].
 Finally, replay does not allow to unlearn specific examples easily, e.g., in case they are not needed anymore.
-Other CL approaches require repeated training from scratch [66], often rendering them infeasible for practical applications.
 We argue that simplicity and flexibility are not necessarily mutually exclusive properties of modern DL approaches.
-In contrast to common, induction-based approaches, where predictions are inferred from general, learned principles, we embrace transduction [21], which is characterized by reasoning from specific training cases to specific test cases.
-Our Work: We propose PARMESAN (parameter-free memory search and transduction), an approach that separates computation from memory and extends transduction to complex, high-dimensional labels for dense prediction tasks while at the same time retaining high flexibility.
+In contrast to common, induction-based approaches, where predictions are inferred from general, learned principles, we embrace transduction, which is characterized by reasoning from specific training cases to specific test cases [19].
+Our Work: We propose PARMESAN (parameter-free memory search and transduction), an approach that separates computation from memory and extends transduction to dense prediction tasks while at the same time retaining high flexibility.
 Our method combines several concepts to perform hierarchical correspondence matching with a memory that contains task-specific, labeled data.
 Rather than requiring to incorporate all knowledge into learnable parameters, the memory in our method extends the capacity of existing models.
-We further propose a message passing approach to take advantage of correlations within a query sample.
-In general, PARMESAN is parameter-free in the sense that it does not have learnable parameters.
-Instead, learning can be easily and intuitively performed by non-ML experts via memory consolidation, i.e., adding, removing, or modifying memory samples.
+We further propose a message passing approach to take advantage of correlations within a query sample, as well as a sparsity approach to retain memory efficiency.
+PARMESAN is parameter-free in the sense that it does not have learnable parameters.
+Instead, learning can be easily performed via memory consolidation, i.e., adding, removing, or modifying memory samples.
 Our method allows for fast learning and unlearning of individual examples or parts thereof.
-Unlike many DL approaches for CL and FSL, our method does not require any continual training or fine-tuning.
-It therefore neither suffers from parameter-induced forgetting and memory overfitting, nor does it require complicated and energy-hungry training by design.
-While we apply our method to 2D images in this work, it canonically transfers to other data dimensionalities such as 3D volumes and 1D sequences, as well as being compatible with spatio-temporal data.
-We do not impose major restrictions on memory size and allow for flexible memory management.
-Our method can be applied in combination with common DL architectures and is task-agnostic, i.e., transfers to different dense prediction tasks as well as multi-tasking.
-Our method can be easily integrated into existing DL architectures, thereby enhancing both flexibility and learning speed.
+Unlike most CL methods, our approach does not require any continual parameter-training by design.
+It therefore neither suffers from parameter-induced forgetting and memory overfitting, nor does it require complicated and energy-hungry training.
+PARMESAN can handle various data dimensionalities such as 1D sequences, 2D images, 3D volumes, as well as spatio-temporal data.
+We do not impose major restrictions on memory size and allow flexible memory management.
+Our method can be applied in combination with common DL architectures, thereby enhancing learning speed.
+This is relevant, e.g., in test-time learning or when dealing with limited computational resources.
+It is also task-agnostic, i.e., it transfers to different dense prediction tasks as well as multi-tasking.
 Our Contributions:
-  • We propose a flexible, parameter-free transduction method for dense prediction tasks using hierarchical correspondence matching with a memory module.
-  • We further propose a parameter-free message passing approach to take advantage of intra-query correlations.
-  • We demonstrate that PARMESAN can be used in combination with common feature extractors.
-  • We show that our method can be successfully applied to CL and FSL, being on par with commonly used baselines in terms of predictive performance.
-At the same time, our method learns up to 370 times faster, has stable knowledge retention properties, and is data-efficient.
-We demonstrate the capabilities of our approach for semantic segmentation and monocular depth estimation on the Cityscapes and JSRT Chest X-ray datasets.
+  • A parameter-free transduction method for dense prediction tasks using hierarchical correspondence matching with a memory.
+  • A parameter-free message passing approach to take advantage of intra-query correlations.
+  • A novelty-sparsity approach to achieve high memory diversity while remaining hardware-efficient.
+  • A demonstration of our method’s flexibility by using it in combination with common DL archictectures for semantic segmentation and depth estimation using the Cityscapes and JSRT Chest X-ray datasets.
+  • We show that our method can be successfully applied to CL, being on par with well-established baselines in terms of predictive performance and memory efficiency while at the same time learning by 3-4 orders of magnitude faster and having stable knowledge retention.
 2 Related work
-Our method combines concepts from the following fields:
-Memory Networks: Recently, the field of memory networks [31, 106, 25, 26, 68] experiences renewed interest.
-When separating computation and memory [26], some parts of an ML model are responsible for computing, others for storing knowledge.
-With this notion it is possible to add new information easily, while at the same time being able to treat memory contents as variables.
+Our method is a novel combination of concepts from the following fields:
+Memory Networks: Recently, memory networks experience renewed interest [27, 94, 22, 23, 62, 55, 1].
+The notion of separating computation and memory [23] allows incorporating new information easily while at the same time enabling treatment of memory contents as variables.
 Memory consolidation [20, 3] concerns the strategy for deciding which knowledge to keep in the presence of limited memory resources.
-Sophisticated memory consolidation methods may include, e.g., novelty search [44], where more informative examples replace less informative examples in memory.
-Transduction: Recently, transduction methods experience a revival due to their effectiveness on various classification tasks [2, 37].
-One of the most prominent transduction methods is based on k-nearest neighbors (k-NN) [16], which aims to find similar examples w.r.t.
+A sophisticated memory consolidation method is novelty search [38], where only the most informative examples are kept.
+We expand on this approach when introducing sparsity to stored samples.
+Transduction: Recently, transduction methods experience a revival due to their effectiveness in computer vision [2].
+A well-established transduction method is k-nearest neighbors ($k$-NN) [14, 32], which aims to find similar examples w.r.t.
 common features.
 Transduction via $k$-NN was demonstrated to work well for image classification [54], where labeled training samples and their hidden representations are stored in memory.
-Given a query sample, labels of similar samples are retrieved from memory, followed by a majority voting for prediction.
-Moreover, k-NN in combination with vision transformers [98, 11] yielded promising results for dense out-of-distribution (OOD) detection [20].
+Moreover, $k$-NN in combination with vision transformers [89, 11] yielded promising results for dense out-of-distribution detection [18].
 However, making $k$-NN scalable to dense prediction tasks at high resolutions remains underexplored.
-Correspondence Matching: Dense correspondence matching [75, 105, 48] [53, 36, 32, 86, 13] is a fundamental problem in computer vision and often acts as a starting point for various downstream tasks such as optical flow, depth-estimation, etc.
-In this context, promising methods for improving feature correlation have been proposed [94, 34].
-Also, methods for exploiting pyramid features have been introduced for object detection [24, 47] and semantic feature matching [95, 111].
-Recently, large-scale foundation models [4] have been applied for one-shot semantic segmentation [49].
-Many existing matching approaches use a single reference sample with known correspondences, but typically do not scale well to more than one reference sample.
-Moreover, most methods rely on specific assumptions commonly found in setups with videos or stereo or multi-view images.
+Correspondence Matching: Dense correspondence matching [67, 93, 43, 48, 31, 28, 78, 12] is a fundamental problem in computer vision, often acting as a starting point for downstream tasks like optical flow and depth-estimation.
+In this context, promising methods to improve feature correlation have been proposed [85, 29, 50].
+Also, feature pyramids were exploited for object detection [21, 42] and semantic feature matching [86, 98].
+Recently, foundation models [4] have been applied for feature matching in one-shot semantic segmentation [44].
+Many existing matching approaches use a single reference sample with known correspondences, but typically can not handle more than one reference sample.
+Moreover, most methods rely on specific assumptions commonly found in setups with videos or stereo / multi-view images.
 In contrast, our proposed method allows matching of arbitrary samples and does not require strong assumptions regarding the presence of correspondences.
-Message Passing (MP): Graph neural networks [81, 9, 40] introduced the concept of message passing to DL, where graph nodes are updated depending on received messages from connected nodes.
+Message Passing (MP): Graph neural networks [74, 10, 35] introduced the concept of MP to DL, where graph nodes are updated depending on received messages from connected nodes.
 MP was successfully applied, e.g., to various problems in bioinformatics, material science, and chemistry [97, 64].
-Continual Learning (CL): The field of CL [93, 73, 92, 61, 8, 102] studies the ability of an ML model to acquire, update, and accumulate knowledge in an incremental manner.
-This enables models, e.g., to adapt to data-distribution shifts and entirely new tasks.
+Continual Learning (CL): The field of CL [84, 65, 83, 56, 9, 91, 96] studies the ability of an ML model to acquire, update, and accumulate knowledge in an incremental manner.
+This enables adaptation to data-distribution shifts [61, 79] and new tasks.
 CL is also relevant when frequent re-training is infeasible.
-Three CL scenarios of increasing difficulty have been proposed [97]: task-incremental (TI), domain-incremental (DI), and class-incremental (CI) CL.
-The definition of the CI scenario is not limited to classification tasks, but is characterized by the ability of a model to infer which task it is presented with.
-Prominent approaches for CL include regularization-based methods [41, 46], parameter-isolation methods [54, 80], and replay methods [82, 71, 51, 76].
-In contrast to biological neural networks, replay in ML is used for regularizing ML models, where the idea is to learn from new data in tandem with data that is stored in a memory or generated during training.
-Replay methods can even lead to optimal CL in case of a perfect memory [42].
-The notion of knowledge retention and mitigation of forgetting [27] plays a major role in learning.
-Forgetting is affected by many aspects such as capacity, plasticity, and the quality as well as diversity of training data.
-The catastrophic forgetting problem in CL [56, 18, 55, 19, 41] refers to learnable parameters adapting to the most recently seen training examples, thereby discarding knowledge from old training examples.
-Few-Shot Learning (FSL): Many DL approaches lack the ability to learn efficiently when data is limited.
-This led to advances in the field of FSL [33], i.e., learning from only a few labeled examples.
-Available examples are also referred to as support samples.
-Some methods approach FSL by the means of meta learning [12, 69, 22], where an outer optimization loop learns a strategy for an inner loop, which further adjusts a model to a specific task.
-The outer loop is performed using large datasets, whereas the inner loop has access to support samples.
-Another approach to FSL is transfer learning via fine-tuning [15, 60, 100, 1, 45].
-Starting from a pre-trained model, learnable parameters or parts thereof are fine-tuned to a specific task using available support samples.
-Initializing models via pre-training on similar data and tasks typically benefits finding models with high performance after fine-tuning.
-Privacy: Common application domains for CL and FSL are autonomous driving and medicine, where data distribution shifts [67] over time are common [87, 64] and data is often scarce.
-However, storing any kind of personally identifiable data such as patient information in a memory can raise privacy concerns.
-Even storing hidden representations [29] or detailed, sample-specific labels, raises data-privacy issues due to the possibility of reconstruction attacks [17, 85].
-The field of data-free CL [46] has emerged as a viable extension when privacy-preservation is required.
+Three CL scenarios of increasing difficulty have been proposed [88]: task-incremental (TI), domain-incremental (DI), and class-incremental (CI).
+The definition of the CI scenario is not limited to classification, but is characterized by the ability of a model to infer which task it is presented with.
+Prominent CL approaches are based on regularization [36, 41], parameter-isolation [49, 72], and replay [75, 63, 46, 68, 37].
+In contrast to biological neural networks, replay in ML is used for regularization, where the idea is to learn from new data in tandem with data that is stored in a memory.
+In CL, the notion of knowledge retention [24] plays a major role in learning.
+Forgetting is affected by aspects such as capacity, plasticity, and the quality as well as diversity of training data.
+The catastrophic forgetting problem in CL [52, 16, 51, 17, 36] refers to learnable parameters adapting to the most recently seen training data, thereby discarding knowledge from old data.
+Note that storing sensitive data such as patient information in a memory raises privacy concerns due to the possibility of reconstruction attacks [15, 77].
+The field of data-free CL [41] using generated [75] or distilled [92, 25] data has emerged as a viable extension when privacy-preservation is needed.
 3 Method
-We now describe our parameter-free transduction method in more detail (also see Fig.
-1).
-First, we provide an overview of our method and describe the individual components of our model, which consists of a frozen feature extractor, a memory, and a transduction module (Sec.
-3.1).
-Given extracted features of a query input, we perform transductive reasoning by searching for similar features within memory samples and retrieving stored labels (Sec.
-3.2).
-Finally, we take advantage of intra-query correlations via message passing (Sec.
-3.3) and further improve predictions by applying test-time augmentation (Sec.
-3.4).
+In this section we describe our parameter-free transduction method (cf.
+Fig.
+1) in more detail.
+First, we provide an overview of our method and describe the individual components of our model, which consists of a frozen feature extractor, a memory, and a transduction module (Section 3.1).
+Given a query input, we perform dense hierarchical correspondence matching with memory contents and take advantage of intra-query correlations via message passing (Section 3.2).
+Finally, we improve memory efficiency by introducing sparsity to stored samples (Section 3.3).
 3.1 General Setup of PARMESAN
-Feature Extraction: Our method requires a pre-trained feature extractor $F$.
-Let $F$ be a neural network with learnable parameters $\theta$ and consisting of an encoder-like or encoder-decoder-like architecture (e.g., U-Net, ConvNeXt-v2, SwinUNETR, etc.) that delivers a feature pyramid $(h)^{l}$ of hidden representations with levels $l=1,...,n$.
-We refer to $h^{1}$ as the hidden representation of the first level and $h^{n}$ as the hidden representation of the final level (bottleneck).
-Individual “pixels” in $(h)^{l}$ with their respective features are referred to as nodes.
+Our proposed method requires a pre-trained feature extractor $F$.
+Let $F$ be a neural network with learnable parameters $\theta$ consisting of an encoder-like or encoder-decoder-like architecture (e.g., [69, 73, 80]) that delivers a feature pyramid $(h)^{l}$ of hidden representations with levels $l=1,...,n$.
+We refer to $h^{1}$ as the hidden representation of $l=1$ (high-res) and $h^{n}$ as the hidden representation of $l=n$ (low-res).
+Individual “pixels” in $(h)^{l}$ with their respective features are called “nodes”.
 The total number of nodes for level $l$ is $p^{l}=\prod_{dim}res_{dim}^{l}$, where $dim$ refers to the data dimension and $res^{l}$ to the grid resolutions.
-Our proposed method is designed with the ability to perform tasks like CL and FSL with frozen parameters $\theta$.
-One can either pre-train $F$ in a supervised or unsupervised manner, or can use existing pre-trained models.
-We recommend pre-training $F$ on a dense prediction task in order to enrich features with dense information.
-Parameter-free Transduction: Let $X$, $H$, and $Z$ be the input-, latent-, and output spaces, respectively.
+Our method is designed to perform CL with frozen parameters $\theta$.
+$F$ can be pre-trained in a supervised or unsupervised manner, or can be taken from public repositories.
+Although not strictly required, we recommend pre-training $F$ on a dense prediction task in order to enrich features with dense information.
+Let $X$, $H$, and $Z$ be the input-, latent-, and output spaces, respectively.
 Densely labeled data is referred to as $(x,z)$ with inputs $x\in X$ and labels $z\in Z$.
 We define a solver $S:X\rightarrow Z$ as the combination of a feature extractor $F:X\rightarrow H$, a memory $M$, and a parameter-free transduction module $G:H\rightarrow Z$.
-Given a query input $x_{q}$ and a memory $M$ that stores $m$ labeled samples, $G$ then predicts $y_{q}=G(F(x_{q};\theta),M)$.
-In essence, $G$ performs hierarchical correspondence matching between a query feature pyramid $(h)_{q}^{l}$ and the memory content, followed by label-retrieval and reduction.
+Given a query input $x_{q}$ and a memory $M$ that stores $m$ labeled samples, $G$ predicts $y_{q}=G(F(x_{q};\theta),M)$.
+$G$ performs hierarchical correspondence matching between a query feature pyramid $(h)_{q}^{l}$ and memory contents, followed by label-retrieval and postprocessing.
 We use $M$ for both learning and long-term knowledge retention.
 Knowledge in $M$ can be modified explicitly, allowing for nuanced and sensible management of what to learn, what to retain, and what to forget.
-Note that this setup shifts the plasticity-robustness trade-off from learnable parameters $\theta$ with parameter-induced forgetting to $M$ with memory-induced forgetting.
-Assumptions: We assume grid-based data structures such as images, volumes, videos, and time series.
+Note that this setup shifts the plasticity-robustness trade-off from learnable parameters with parameter-induced forgetting to a memory with memory-induced forgetting.
+In contrast to other approaches, our method is designed to learn fast in the sense that it does not require adapting $\theta$ to perform complex tasks such as CL.
+Our method only requires memory consolidation, which in our case means saving examples in $M$ and extracting their feature pyramids $(h)_{j}^{l}$ in a single forward pass with frozen $\theta$.
+We assume grid-based data structures such as images, volumes, and time series.
 Although we focus on orthogonal grids in this work, our method can be transferred to irregular, adaptive, and isometric grids.
@@
 When taken to extremes, individual tokens of labels can even represent entire data structures such as sets, series, or graphs.
-Privacy Preservation: Our method also extends to privacy-preserving settings, where privacy leaks can be mitigated by storing generated [82] or distilled [103, 28] data in $M$.
-We do not impose strict requirements for having knowledge about labels, e.g., class segments, and their distribution, e.g., class abundances.
-This setup even allows for keeping memory labels private if needed.
-3.2 Hierarchical Correspondence Matching
-We initialize our model by filling $M$ with $m$ labeled samples consisting of $(x,z)_{j}$ as well as their respective feature pyramids $(h)_{j}^{l}$, where $j=1,...,m$ and $l=1,...,n$.
-Leaf nodes in level $l=0$ refer to memory labels $z_{j}$.
+3.2 Correspondence Matching and Intra-Query Transduction
+We initialize our model by filling $M$ with $m$ labeled samples $(x,z,(h)^{l})_{j}$, where $j=1,...,m$ and $l=1,...,n$.
+Leaf nodes in $l=0$ refer to $z_{j}$.
 Next, we define connectivity kernels which define children-parent relations across different levels.
-Kernels remain fixed for every sample and for every level.
-For more details, see Appendix 0.A.
-Some neural architectures may not provide desired feature pyramids in the sense that resolutions are halved at every level or that bottlenecks have small resolutions.
-In these cases, connectivity kernels can either be adjusted accordingly, or missing representations can be inserted by interpolation or extrapolation of neighboring representations.
+Kernels remain fixed for every sample and every level (also see Appendix 0.A).
 Algorithm 1 PARMESAN model inference
@@
 return $y_{q,raw}$triangleright raw prediction
-We formulate the transductive forward pass of our method in Algorithm 1.
-First, a query $x_{q}$ is encoded via $F$ to obtain $(h)_{q}^{l}$.
-Then, $G$ takes $(h)_{q}^{l}$ together with memory feature pyramids $(h)_{j}^{l}$ and their respective labels $z_{j}$ as inputs.
-The overall objective is searching for both globally and locally similar nodes in $M$.
-Starting from $l=n$, the search objective in each level is first comparing queries $h_{q}^{l}$ with keys $h_{j}^{l}$, followed by retrieving the children of the top-$k$ most similar memory nodes.
-We use the cosine similarity between nodes’ hidden features as our similarity measure.
-The accumulated similarity $s_{acc}^{l}$ between two nodes at level $l$ is defined as the product of their similarities $s^{l}$ at level $l$ and the similarities $s_{k,par}$ of their respective parents across all higher levels:
-$s_{acc}^{l}=\prod_{u=l}^{n}sim(h_{q}^{u},h_{j}^{u})=\prod_{u=l}^{n}\frac{h_{q}^{u}\cdot h_{j}^{u}}{\lVert h_{q}^{u}\rVert\ \lVert h_{j}^{u}\rVert}$ (1)
-For each of the $p^{l}$ query nodes, we use $s_{acc}^{l}$ to identify the top-$k$ most similar nodes in level $l$ and retrieve their respective children in level $l-1$.
-This implies that the top-$k$ most similar nodes in level $l$ become parent nodes for level $l-1$.
-Children nodes in level $l-1$ are required for processing level $l-1$, thereby forming a hierarchy across levels.
-The hyperparameter $\phi\in(0.0,1.0]$ controls the number $k$ of top-$k$ nodes to retain for every query node in each level during search.
-Effectively, $\phi$ is the rate of reduction for $k$ in each level.
+We formalize the forward pass of our method in Algorithm 1.
+First, a query $x_{q}$ is encoded by $F$ to get $(h)_{q}^{l}$.
+Then, $G$ takes $(h)_{q}^{l}$ together with stored $(h)_{j}^{l}$ and $z_{j}$ as inputs.
+The overall objective is to find globally and locally similar patterns in $M$ w.r.t.
+$x_{q}$.
+Traversing levels backwards starting from $l=n$, we compare queries $h_{q}^{l}$ with keys $h_{j}^{l}$, followed by retrieving the children (within $l-1$) of the top-$k$ most similar nodes of $h_{j}^{l}$.
+We use cosine similarity by default, where the accumulated similarity $s_{acc}^{l}$ between two nodes at level $l$ is the product of their similarities $s^{l}$ at level $l$ and the similarities $s_{k,par}$ of their respective parents across all higher levels:
+$s_{acc}^{l}=\prod_{u=l}^{n}sim(h_{q}^{u},h_{j}^{u})=\prod_{u=l}^{n}\frac{h_{q}^{u}\cdot h_{j}^{u}}{\lVert h_{q}^{u}\rVert\ \lVert h_{j}^{u}\rVert}\ .$ (1)
+For each of the $p^{l}$ query nodes, we use $s_{acc}^{l}$ for comparisons.
+The top-$k$ most similar nodes in $l$ become parents for $l-1$.
+Children in $l-1$ are required for processing $l-1$, thereby forming a hierarchy across levels.
+The reduction rate $\phi\in(0.0,1.0]$ controls $k$, i.e., how many nodes are retained in each level.
 A variable number of nodes can be retained in different memory samples.
-Considering the data dimension $dim$ and the grid resolutions $res^{l}$, exhaustive search would result in a computational complexity of $\mathcal{O}((\prod_{dim}res_{dim}^{l})^{2})$ for level $l$.
-In contrast, our approach performs search locally and has a computational complexity of only $\mathcal{O}(n_{ch}\cdot(\prod_{dim}res_{dim}^{l}))$ with $n_{ch}$ being the number of child nodes in level $l+1$.
-Also, since search is performed independently for every query node, our approach can be highly parallelized in practice and is thus scalable to high resolutions and high-dimensional data.
-Moreover, since $h^{l}$ are strongly correlated across levels, there is a high chance that similarities to query nodes are also strongly correlated across levels.
-In other words, searching locally pre-filters nodes that are unlikely to be among the top-$k$ most similar nodes as search progresses to the next level.
 Memory search results in a dense correspondence map between a query $x_{q}$ and, for every query node, their top-$k$ most similar memory nodes $i_{k,par}$.
 We use these matches to retrieve task-specific labels $z_{k}$ from $M$.
-To obtain a raw prediction $y_{q,raw}$, we use the softmax-normalized similarities $s_{k,par}$ of all matches to obtain a weighted average of retrieved labels.
-3.3 Intra-Query Transduction via Message Passing
-Up until this point we did not exploit any intra-query correlations such as applying geometric priors or enforcing smoothness of predictions.
+We obtain a raw prediction by weighting retrieved labels with softmax-normalized similarities $s_{k,par}$ of all matches.
+Considering the data dimension $dim$ and the grid resolutions $res^{l}$, exhaustive search would result in a computational complexity of $\mathcal{O}((\prod_{dim}res_{dim}^{l})^{2})$ for level $l$.
+In contrast, our method performs local search and has a complexity of only $\mathcal{O}(n_{ch}\cdot(\prod_{dim}res_{dim}^{l}))$ with $n_{ch}$ being the number of children of nodes in $l+1$.
+Search is performed independently and in parallel for every query node, making our approach scalable to high resolutions.
+Moreover, since $h^{l}$ are strongly correlated across levels, there is a high chance that similarities to query nodes are also strongly correlated across levels.
+In other words, searching locally pre-filters nodes that are unlikely to be among the top-$k$ most similar nodes in the next level anyway.
+Up until this point, no intra-query correlations had been exploited by using geometric or smoothness priors.
 Strong local correlations are common in grid-based data and should be exploited if possible.
-The simplest variant of intra-query transduction is Gaussian blurring, where local smoothness over neighboring nodes is enforced.
-However, blurring is not content-aware and can harm performance at certain regions, e.g., at sharp corners or small objects.
 We therefore propose a content-aware, parameter-free message passing (MP) approach to account for local correlations within the query and achieve pixel-exact predictions.
-We re-use our hierarchical search without memory content to find the top-$\kappa$ nearest neighbor nodes within the query sample itself.
+Re-using our hierarchical search without memory content, top-$\kappa$ nearest neighbor nodes within the query itself are found.
 Similarity scores between a query node $i$ and its nearest neighbors $j$ are referred to as edges $e_{ij}$.
-We then perform MP to refine raw predictions $y_{q,raw}$.
+MP is then perfomed to refine raw predictions $y_{q,raw}$.
 Considering a query node state $y_{q,i}$, we define the equations for the edges $e_{ij}$, the aggregated message $\hat{y}_{q,i}$, and the node update from step $t\rightarrow t+1$ as
@@
 The hyperparameter $\lambda\in(0.0,1.0]$ controls the overall strength of MP per step.
-We apply Eq.
-3 and Eq.
-4 repeatedly for all query nodes in level $l=0$ until convergence [113], i.e., when $||y_{q,i}^{t+1}-{y}_{q,i}^{t}||\rightarrow 0$.
-The resulting prediction $y_{q}$ then contains less artifacts and is spatially more consistent compared to $y_{q,raw}$.
-3.4 Test-Time Augmentation
-Similar to recent work [36], we perform test-time augmentation (TTA) to resolve small errors.
-We apply multiple forward passes with down-scaled and reflection-padded variants of $x_{q}$.
-Registration of down-scaled variants with the original variant is achieved by up-scaling and center-cropping the respective predictions.
-We obtain a final prediction $y_{q}$ by averaging over all passes.
-3.5 Fast Learning
-PARMESAN is designed to learn fast in the sense that it does not require training of learnable parameters $\theta$ to perform complex tasks such as CL or FSL.
-In contrast to other approaches, our method only requires memory consolidation, i.e., saving examples in $M$ and extracting their respective feature pyramids $(h)_{j}^{l}$ in a single forward pass using $F$ with pre-trained, frozen $\theta$.
-For a given hardware setup, learning speed can be split up into required wall-clock times $\tau_{p}$ and $\tau_{l}$ for pre-training and learning, respectively.
-We define $\tau_{p}$ on a per-dataset basis and $\tau_{l}$ on a per-sample basis.
-The total required wall-clock time to learn $n$ training samples is then $T=\tau_{p}+n\cdot\tau_{l}$.
+We apply Equations 3 and 4 repeatedly for all query nodes in level $l=0$ until convergence [100], i.e., when $||y_{q,i}^{t+1}-{y}_{q,i}^{t}||\rightarrow 0$.
+The resulting prediction $y_{q}$ then contains fewer artifacts and is spatially more consistent compared to $y_{q,raw}$.
+3.3 Memory-Efficiency via Sparsity and Novelty
+In its basic setup, PARMESAN requires considerable hardware memory resources to operate.
+We thus introduce sparsity, i.e., a reduction of the number of patterns per stored sample, to become exponentially more memory efficient (see Fig.
+2).
+Sparsity effectively reduces spatial resolutions of stored samples and is applied to both $(h)_{j}^{l}$ and $z_{j}$ after feature extraction of full-resolution inputs $x_{j}$.
+Starting from $l=1$, the number of sparse levels $n_{sp}$ defines the overall degree of sparsity.
+Naive sparsity can be achieved by random subsampling of nodes in levels $l=1,...,n_{sp}$.
+However, we apply local novelty search to retain high diversity in memory contents.
+Starting from $l=1$, we keep the most novel node in every $2^{dim}$ patch such that sparsified resolution $res_{sp,dim}^{1}=res_{dim}^{2}$.
+After applying sparsity to $l=2$ such that $res_{sp,dim}^{2}=res_{dim}^{3}$, we discard previously kept nodes in $l=1$ such that $res_{sp,dim}^{1}=res_{sp,dim}^{2}=res_{dim}^{3}$.
+We repeat this procedure for every sparse level and independently for all memory samples.
+Figure 2: Memory sample feature pyramid with $n_{sp}=2$ sparse levels obtained via iterative, local novelty search (indicated by circles). Filled nodes are kept in $M$.
 4 Experiments and Results
-We demonstrate the capabilities of our method in three groups of experiments:
-A0 – A18:
-ablation studies for semantic segmentation and monocular depth estimation to demonstrate flexibility (Sec. 4.1, Tab. 1, and Tab. 2),
-B0 – B7:
-CL experiments for studying knowledge retention properties of our method (Sec. 4.2, Tab. 3, and Tab. 4),
-C0 – C7:
-FSL experiments to investigate data-efficiency (Sec. 4.3 and Tab. 5).
-More details regarding applied models, training and evaluation, as well as qualitative results for all experiments are provided in Sec. 0.B.2.
-Our main goal is to improve flexibility and adaptation capabilities of DL models while simplifying and speeding up learning for CL and FSL.
-Problem-specific methods are expected to outperform more general methods such as ours [57, 104, 101].
-Datasets: We use the Cityscapes (CITY) benchmark dataset [7] for semantic segmentation and monocular depth estimation.
-Semantic segmentation masks comprise fine and coarse labels, whereas disparity maps are provided for depth estimation.
+We demonstrate the capabilities of our method in ablation studies (Section 4.1) and CL experiments (Section 4.2).
+For ablations, we aim to demonstrate flexibility and efficiency while simplifying and speeding up learning.
+For CL, we study knowledge retention and overall performance.
+More details regarding applied models, training and evaluation, and additional results are provided in Section 0.B.1 and 0.B.2.
+Datasets: We use the Cityscapes (CITY) dataset [8] for semantic segmentation and monocular depth estimation.
+Segmentation masks comprise fine and coarse labels.
+Disparity maps are given for depth estimation.
 We also use the JSRT Chest X-ray dataset [76] for semantic segmentation.
-We refer to Sec.
-0.B.1 for further details.
-Method: We employ a U-Net [77] variant as our default DL model and feature extractor.
-For experiments A11 and A14 we use ConvNeXt Tiny [50] with $\theta$ trained on ADE20K [112].
-For all experiments that require $M$, we select memory samples randomly from the training split while keeping the random generation process fixed to ensure comparability.
-For PARMESAN, we set $m~{}=~{}256$, $\phi=0.5$, apply message passing (MP) and test-time augmentation (TTA), and use both pre-trained U-Net encoder and decoder for feature extraction unless stated otherwise.
-We use Gaussian blurring with $\sigma=5$ pixels in experiment A2.
-CL and FSL are performed by memory consolidation of training samples.
-To obtain memory samples for the privacy-preserving experiment A6, we additionally apply data distillation via linear mixup of two random training samples [109, 43, 83, 70].
-Baselines: Joint training (JOINT) refers to training on all available training data, thus representing the default supervised learning setup and the upper bound for CL [97].
-Fine-tuning (FT) is a simple, yet strong baseline for FSL [10].
-Due to catastrophic forgetting, however, FT is regarded as the lower bound for CL [97].
-We further employ memory-based approaches such as classical replay (REPLAY) [82, 71, 51, 76] and greedy-sampling dumb learning (GDUMB) [66] as additional baselines for CL.
-With GDUMB, training is performed from scratch for each individual CL step, thereby exclusively using memory samples.
-For consistent and fair comparisons to our method, we implement all baselines and adjust method-specific hyperparameters to our data setup.
+Method: We employ a U-Net variant as our default DL model and feature extractor $F$.
+Some experiments use a ConvNeXt Tiny [45] encoder pre-trained on ADE20K [99].
+If $M$ is used, we select memory samples randomly from the training split while keeping the random generation process fixed across all experiments.
+For PARMESAN, we set $m\!=\!2975$, $\phi\!=\!0.5$, $n_{sp}\!=\!3$ ($m\!=\!199$, $\phi\!=\!0.8$, $n_{sp}\!=\!3$ for JSRT), apply message passing (MP), and take decoder features from $F$ unless stated otherwise.
+CL is performed by memory consolidation of training samples.
+We also use test-time augmentation (TTA) to boost performance [31].
+Baselines: We select baselines with the premise to compare to representative and well-established methods in the field.
+Joint training (JOINT) refers to training on all training data, thus representing the default supervised learning setup and the upper bound for CL [88].
+Due to catastrophic forgetting, fine-tuning (FT) is regarded as the lower bound for CL [88].
+We employ classical replay (REPLAY) [75, 63, 46, 68], greedy-sampling dumb learning (GDUMB) [60], and Modeling the Background (MiB) [5] as additional CL baselines.
+Training via GDUMB is performed from scratch for each individual CL step, thereby exclusively using memory samples.
+For fair and consistent comparisons, we implement all baselines and adjust method-specific hyperparameters to our data setup.
 Evaluation: We use the mean Intersection over Union of classes ($\text{mIoU}_{\text{cl}}$) and categories ($\text{mIoU}_{\text{cat}}$) as evaluation metrics for semantic segmentation and Root-Mean-Square Error (RMSE) in meters for depth estimation.
 We report best performances on the CITY validation set and on the JSRT test set, since labels for the CITY test set are not publicly available.
+For a given hardware setup, the learning speed $\tau_{l}$ is the average wall-clock time required to learn a training sample.
+The total learning time for $n$ samples is then $T=n\cdot\tau_{l}$.
 4.1 Ablation Studies
-We perform ablation studies of various components to analyze the impact on learning speed and predictive performance, as well as demonstrating the flexibility of our method with different architectures of $F$.
-We summarize our results in Tab.
-1 (segmentation) and Tab.
-2 (depth) and visualize predictions in Fig.
-2 (CITY, segmentation), Fig.
-3 (CITY, depth), and Fig.
-4 (JSRT, segmentation).
-Learning with our method can be done in less than 0.1 seconds per sample (see column $\tau_{l}$ in Tab.
-1 and Tab.
-2), thereby mitigating the demand for additional computational costs when attaching PARMESAN to existing models.
-Our method outperforms JOINT on CITY when trained on coarse labels (A9, A10) and evaluated on fine labels.
-As expected, our method under-performs JOINT when using fine labels (A0, A1), on JSRT (A12, A13, A14), and for depth estimation (A15, A16).
-Using only the encoder (A7) or using an entirely untrained, randomly initialized model (A8) yields promising results already, indicating that our transduction module $G$ imposes useful inductive biases.
-PARMESAN can be successfully used in combination with a ConvNeXt feature extractor on both CITY and JSRT when trained on ADE20K [50] (A11, A14).
-MP and TTA (A1, A3, A4, A16, A17) considerably improve performance over raw predictions (A5, A18).
-As expected, content-aware MP (A1) also outperforms non-content-aware Gaussian blurring (A2).
-Comparing panels A1 and A5 in Fig.
-2, MP and TTA succeed in reducing small artifacts while retaining sharp corners of predicted segments.
-In contrast, blurring typically comes at the cost of missing small segments and sharp corners.
-Privacy-preservation by data distillation (A6) slightly decreases performance compared to the non-private setup (A1).
+We perform ablation studies of various components to analyze the impact on learning speed and predictive performance, as well as demonstrating the flexibility of our method with different architectures of $F$ (see Table 1).
+Performance of our method is on-par with JOINT (A0 / A1, A9 / A10, A12 / A13) while greatly enhancing flexibility and learning speed.
+Learning with PARMESAN can be done in $\tau_{l}\simeq 0.01$ seconds per sample, thereby mitigating the demand for additional computational costs when attached to existing models.
+Using only the encoder (A4) or even random weights (A5) yields promising results already, indicating that $G$ imposes useful inductive biases.
+PARMESAN can be successfully used for transfer learning when using a publicly available ConvNeXt pre-trained on ADE20K (A8, A11) or pre-training on coarse labels (A7).
+MP and TTA (A1 / A2 / A3) considerably improve raw predictions and succeed in reducing small artifacts while retaining sharp corners of predicted segments (see Fig.
+3).
 Figure 3: Model predictions and analysis of PARMESAN.
@@
 The idx panel visualizes indices of retrieved nearest neighbor nodes from $M$.
-Similar colors indicate that nodes are retrieved from close-by memory locations, most likely from within the same samples.
-The sim panel visualizes the respective similarity scores, where bright regions refer to high similarities.
+Similar colors indicate labels retrieved from close-by memory locations, most likely from within same samples.
+The sim panel visualizes $s_{acc}^{1}$, where bright regions refer to high similarities.
 We also show 3 query pixels and their matches in $M$, indicating both global and local semantic correspondence.
-Right: PARMESAN can be applied flexibly (A5, A11) and perform complex tasks like CL (B5, B7) and FSL (C3, C7) via memory consolidation.
-Figure 3: Model predictions for monocular depth estimation experiments.
-Figure 4: Model predictions for semantic segmentation experiments using JSRT data.
+Right: Flexible use of PARMESAN in various setups.
+Bottom: Depth estimation on CITY and semantic segmentation on JSRT.
 In general, we observe global and local semantic correspondence of retrieved nearest neighbors from $M$, allowing for future work on various topics related to interpretability and explainability of predictions.
@@
 As visible in the sim panel in Fig.
-2, similarity scores seem to be correlated with the heterogeneity of regions in the image.
-Table 1: Results for ablation studies showing semantic segmentation performance for classes ($\text{mIoU}_{\text{cl}}$) and categories ($\text{mIoU}_{\text{cat}}$).
-Column “$F$ pre-train/train“ indicates if the method uses a pre-trained feature extractor $F$ and if any additional training of learnable parameters $\theta$ is required.
-$\tau_{p}$ and $\tau_{l}$ are wall-clock times for pre-training and learning, respectively.
-Experiment IDs marked with * use ConvNeXt pre-trained on ADE20K [50] for feature extraction.
-ID Train data (#samples) Method & Setup $F$ pre-train/train $\tau_{p}$ [h] $\tau_{l}$ [s] $\text{mIoU}_{\text{cl}}$ $\text{mIoU}_{\text{cat}}$
-A0 CITY, fine (2975) JOINT no / yes 0 3.72 82.5 87.4
-A1 CITY, fine (256) ours A0 / no 3.08 0.09 80.0 85.5
-A2 CITY, fine (256) ours, no MP, blur A0 / no 3.08 0.09 79.3 84.8
-A3 CITY, fine (256) ours, no TTA A0 / no 3.08 0.09 77.7 83.5
-A4 CITY, fine (256) ours, no MP A0 / no 3.08 0.09 77.9 83.2
-A5 CITY, fine (256) ours, no MP, no TTA A0 / no 3.08 0.09 75.7 81.6
-A6 CITY, fine (256) ours, private A0 / no 3.08 0.10 77.6 83.2
-A7 CITY, fine (256) ours A0 (E only) / no 3.08 0.08 72.8 79.1
-A8 CITY, fine (256) ours no / no 0 0.09 56.1 63.5
-A9 CITY, coarse (19998) JOINT no / yes 0 0.43 49.6 50.6
-A10 CITY, coarse (256) ours A9 / no 2.41 0.09 50.7 52.5
-A11* CITY, fine (256) ours E only / no - 0.09 74.5 81.4
-A12 JSRT (199) JOINT no / yes 0 23.39 93.3 -
-A13 JSRT (199) ours, $m\!=\!199$, $\phi\!=\!0.8$ A12 / no 1.29 0.02 91.9 -
-A14* JSRT (199) ours, $m\!=\!199$, $\phi\!=\!0.8$ E only / no - 0.02 87.2 -
-Table 2: Results for ablation studies showing depth estimation performance.
-ID Train data (#samples) Method & Setup $F$ pre-train/train $\tau_{p}$ [h] $\tau_{l}$ [s] RMSE [m]
-A15 CITY, disparity (2975) JOINT no / yes 0 6.71 10.5
-A16 CITY, disparity (256) ours A15 / no 5.55 0.03 11.9
-A17 CITY, disparity (256) ours, no MP A15 / no 5.55 0.03 12.5
-A18 CITY, disparity (256) ours, no MP, no TTA A15 / no 5.55 0.03 13.5
+3, similarity scores seem to be correlated with the heterogeneity of regions in the image.
+In Fig. 4, we demonstrate that our novelty-sparsity approach leads to exponentially lower memory requirements while achieving robust predictive performance.
+Table 1: Results for ablation studies, demonstrating the capabilities of PARMESAN.
+For semantic segmentation, we report performance for classes ($\text{mIoU}_{\text{cl}}$) and categories ($\text{mIoU}_{\text{cat}}$).
+“$\theta$ pre-train” and “$\theta$ train” indicate the usage of a pre-trained feature extractor $F$ and if any additional training of learnable parameters $\theta$ is required.
+$\tau_{l}$ is the average wall-clock time for learning a sample.
+ID Data (#train-samples) Method & Setup $\theta$ pre-train $\theta$ train $\tau_{l}$ [s] $\text{mIoU}_{\text{cl}}$ $\text{mIoU}_{\text{cat}}$
+A0 CITY, fine (2975) JOINT 3.72 82.5 87.4
+A1 CITY, fine (2975) ours A0 0.012 82.1 87.0
+A2 CITY, fine (2975) ours, no MP A0 0.012 81.4 86.4
+A3 CITY, fine (2975) ours, no TTA A0 0.012 81.2 86.3
+A4 CITY, fine (2975) ours, E only A0 0.009 75.6 81.6
+A5 CITY, fine (2975) ours, random $\theta$ 0.012 59.6 66.3
+A6 CITY, coarse (19998) JOINT 0.43 49.0 50.3
+A7 CITY, fine (2975) ours A6 0.012 76.9 82.4
+A8 CITY, fine (2975) ours, ConvNeXt (E) [45] 0.021 78.2 84.3
+A9 JSRT (199) JOINT 23.4 93.9 -
+A10 JSRT (199) ours A9 0.009 92.6 -
+A11 JSRT (199) ours, ConvNeXt (E) [45] 0.011 84.7 -
+ID Data (#train-samples) Method & Setup $\theta$ pre-train $\theta$ train $\tau_{l}$ [s] RMSE [m]
+A12 CITY, disparity (2975) JOINT 6.71 10.5
+A13 CITY, disparity (2975) ours A12 0.01 11.5
+Figure 4: Performance w.r.t.
+memory size on CITY.
+We control memory size by introducing sparse levels, leading to exponentially lower memory requirements and stable performance.
+Relevant contents are retained using local novelty search.
 4.2 Continual Learning
-[…]
-Table 3: Results for CL experiments. Performance is measured on all classes after the final CL step (i.e., after 4 steps). Our method is on-par with CL baselines while learning substantially faster.
-ID Train data (#samples) Method & Setup $F$ pre-train/train $\tau_{p}$ [h] $\tau_{l}$ [s] $\text{mIoU}_{\text{cl}}$ $\text{mIoU}_{\text{cat}}$
-B0 CITY, fine (2975) JOINT (upper bound) no / yes 0 20.62 81.7 86.6
-B1 CITY, fine (2975) FT (lower bound) no / yes 0 12.76 4.0 4.3
-B2 CITY, fine (2975) REPLAY, $m\!=\!256$ no / yes 0 5.11 56.2 64.2
-B3 CITY, fine (256) GDUMB, $m\!=\!256$ no / yes 0 110.9 74.4 80.3
-B4 CITY, fine (256) ours B0 ($1^{\text{st}}$ step) / no 3.07 0.32 66.0 70.7
-B5 CITY, fine (256) ours B3 ($1^{\text{st}}$ step) / no 1.42 0.32 65.4 71.0
-B6 CITY, fine (256) GDUMB, $m\!=\!256$ A9 / yes 2.41 63.13 76.9 82.1
-B7 CITY, fine (256) ours A9 / no 2.41 0.32 75.4 81.3
-[…]
-REPLAY (B2) GDUMB (B3) ours (B4) ours (B5) GDUMB (B6) ours (B7)
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-4.3 Few-Shot Learning
-For FSL, we pre-train $F$ on coarse labels (A9).
-We then perform experiments with 10, 8, 6, 4, and 2 fine-labeled support samples.
-FSL via FT is performed by fine-tuning $\theta$ on support samples.
-To study FT under realistic conditions, we divide the support sets into 2 splits with the latter containing 4 samples used for early stopping.
-For FSL with PARMESAN, we fill $M$ with all support samples.
-A low number of support samples can lead to a high variance in performance.
-We thus repeat each experiment 5 times with randomly drawn support samples, where we fix the random generation process across experiments to ensure comparability.
-We report average performances on the full validation set in Tab.
-5.
-Table 5: Results for FSL experiments, averaged over 5 repetitions.
-We use a feature extractor pre-trained on coarse labels and a support set containing fine labels.
-Our method learns fast and is data-efficient especially for small support sets.
-ID Train data (#samples) Method & Setup $F$ pre-train/train $\tau_{p}$ [h] $\tau_{l}$ [s] $\text{mIoU}_{\text{cl}}$ $\text{mIoU}_{\text{cat}}$
-C0 CITY, fine (10, train on 6) FT A9 / yes 2.41 32.29 71.7 76.4
-C1 CITY, fine (8, train on 4) FT A9 / yes 2.41 69.53 71.1 75.9
-C2 CITY, fine (6, train on 2) FT A9 / yes 2.41 225.6 69.5 74.7
-C3 CITY, fine (10) ours, $m\!=\!10$, $\phi\!=\!1$ A9 / no 2.41 0.17 72.2 78.2
-C4 CITY, fine (8) ours, $m\!=\!8$, $\phi\!=\!1$ A9 / no 2.41 0.21 71.1 77.5
-C5 CITY, fine (6) ours, $m\!=\!6$, $\phi\!=\!1$ A9 / no 2.41 0.26 69.9 76.2
-C6 CITY, fine (4) ours, $m\!=\!4$, $\phi\!=\!1$ A9 / no 2.41 0.34 66.9 73.3
-C7 CITY, fine (2) ours, $m\!=\!2$, $\phi\!=\!1$ A9 / no 2.41 0.61 58.3 65.6
-FSL with PARMESAN can be done up to 370 times faster w.r.t.
-FT (C2, C7) and remains on par in terms of predictive performance for equal support set sizes (C0-C5).
-Our method is data-efficient in the sense that is able to utilize the entire support set at inference and that it can be applied in combination with a small support set (C6, C7).
-4.4 Limitations
-Depending on the choice of hyperparameters such as $m$ and $\phi$, PARMESAN might require more computational resources and longer inference times compared to learned task-heads.
-Memory consumption and inference speed can potentially be further improved by introducing sparsity to stored feature pyramids, i.e., reducing the number of unique patterns per memory sample.
-Moreover, choosing a suitable feature extractor might require taking a trade-off between a general-purpose model versus a data- and task-specific model.
+We perform CI-CL experiments, where the goal is to incrementally learn new semantic classes while at the same time automatically inferring the given task.
+We employ two different CL scenarios: 2-2 (4) cat and 13-1 (7) cl [95].
+At 2-2 (4) cat, data at every CL step comprises labels for all classes of different categories: $\textit{D}_{1}$: void & flat, $\textit{D}_{2}$: construction & object, $\textit{D}_{3}$: nature & sky, and $\textit{D}_{4}$: human & vehicle.
+For memory-based methods REPLAY, GDUMB, and PARMESAN, we update samples in $M$ with labels from the current CL step while keeping the stored inputs.
+REPLAY and GDUMB utilize $M$ for training $\theta$, whereas PARMESAN uses $M$ for transduction.
+We summarize our results in Table 2, Table 3, Fig.
+3, and provide additional information in Section 0.B.3.
+FT (B1, B12) and JOINT (B0, B11) behave as expected with strong forgetting and superior performance w.r.t.
+other methods, respectively.
+For same memory sizes, PARMESAN is on par with other CL baselines while enhancing flexibility, allowing simple learning and unlearning of individual examples, and boosting learning speed between 3-4 orders of magnitude.
+We further demonstrate that our method has stable knowledge retention properties (see Table 3).
+This effect can be attributed to our shift towards memory-induced forgetting, where we explicitly state what to learn (new classes), what to retain (old classes), and what to forget (background).
+Table 2: Results for CL experiments on CITY fine, including size of $M$ and learning times $\tau_{l}$. Performance is measured on all classes after the final CL step.
+ID scenario Method & Setup $M$ [GB] $\theta$ pre-train $\theta$ train $\tau_{l}$ [s] $\text{mIoU}_{\text{cl}}$ $\text{mIoU}_{\text{cat}}$
+B0 2-2 (4) cat JOINT - A6 3.64 81.6 86.6
+B1 2-2 (4) cat FT - A6 3.44 4.1 4.3
+B2 2-2 (4) cat REPLAY, $m\!=\!940$ 6.43 A6 2.79 76.8 82.6
+B3 2-2 (4) cat GDUMB, $m\!=\!940$ 6.43 A6 10.7 78.9 84.5
+B4 2-2 (4) cat MiB - A6 6.10 78.8 83.9
+B5 2-2 (4) cat ours, $m\!=\!256$ 0.55 A6 0.012 73.9 79.9
+B6 2-2 (4) cat ours, $m\!=\!940$ 2.02 A6 0.010 75.6 81.3
+B7 2-2 (4) cat ours 6.40 A6 0.012 76.9 82.4
+B8 2-2 (4) cat ours, $m\!=\!256$, ConvNeXt 0.55 [45] 0.021 71.6 78.3
+B9 2-2 (4) cat ours, $m\!=\!940$, ConvNeXt 2.0 [45] 0.019 73.3 80.0
+B10 2-2 (4) cat ours, ConvNeXt 6.35 [45] 0.021 78.2 84.3
+B11 13-1 (7) cl JOINT - A6 3.97 80.4 85.2
+B12 13-1 (7) cl FT - A6 1.89 0.21 0.21
+B13 13-1 (7) cl REPLAY, $m\!=\!940$ 6.43 A6 7.72 79.1 84.1
+B14 13-1 (7) cl GDUMB, $m\!=\!940$ 6.43 A6 9.36 79.1 84.1
+B15 13-1 (7) cl MiB - A6 2.44 65.2 72.5
+B16 13-1 (7) cl ours 6.40 A6 0.001 76.8 82.2
+B17 13-1 (7) cl ours, ConvNeXt 6.35 [45] 0.003 78.1 84.0
+Table 3: Knowledge retention after 4 CL steps, showing absolute changes $\delta$ between initial and final $\text{mIoU}_{\text{cl}}$.
+Datasets $\textit{D}_{i}$ comprise semantic classes of different categories according to the respective CL step.
+Our method has stable knowledge retention and can not suffer from parameter-induced forgetting by design.
+REPLAY (B2) GDUMB (B3) MiB (B4) ours (B6) ours (B7) ours (B9) ours (B10)
+$\textit{D}_{i}$ initial $\delta$ initial $\delta$ initial $\delta$ initial $\delta$ initial $\delta$ initial $\delta$ initial $\delta$
+$\textit{D}_{1}$ 78.70 -2.80 77.89 -0.71 79.73 -2.60 71.21 -0.79 72.05 -1.03 64.39 -0.95 67.99 -0.45
+$\textit{D}_{2}$ 49.66 +3.62 55.87 +0.07 57.21 -3.29 49.05 -0.43 50.52 -0.51 50.17 -0.50 54.40 -0.43
+$\textit{D}_{3}$ 63.05 +4.24 69.78 -0.38 69.12 -1.65 61.68 -0.25 63.12 -0.26 59.27 -0.30 62.47 -0.26
+$\textit{D}_{4}$ 39.60 - 48.37 - 46.67 - 40.33 - 42.03 - 49.27 - 54.68 -
+$avg$ 57.75 +1.69 62.98 -0.25 63.18 -2.51 55.57 -0.49 56.93 -0.60 55.78 -0.58 59.88 -0.38
+$std$ 16.95 3.90 13.32 0.39 14.35 0.82 13.62 0.27 13.29 0.39 7.31 0.33 6.57 0.10
 5 Conclusion
-In this work we proposed PARMESAN, a flexible, parameter-free approach that separates computation from memory and leverages transduction for solving dense prediction tasks.
-We also proposed a message passing approach to take advantage of intra-query correlations.
-Our method can perform complex tasks such as CL and FSL in an easy, intuitive, and optionally private manner without any continuous training of learnable parameters.
-We demonstrated that our approach learns substantially faster than common baselines while achieving competitive predictive performance, showing stable knowledge retention, and being data-efficient.
-Future Work: Possible future work includes leveraging self-supervised feature extractors, studying OOD detection capabilities, as well as investigating memory-related topics such as dataset distillation.
-Performance might be further increased by implementing algorithmic or methodical improvements.
-This includes, e.g., using explorative memory search instead of our current greedy approach, as well as applying novelty search instead of random sampling for memory consolidation.
-Finally, PARMESAN might be beneficial to other fields such as reinforcement learning, where fast adaptation can be important.
+Summary: We propose PARMESAN, a flexible, parameter-free transduction approach for dense prediction tasks.
+We also propose a message passing approach to take advantage of intra-query correlations, as well as a novelty-sparsity approach to retain memory efficiency and diversity.
+Our method is capable of performing CL in an easy and intuitive manner without any continuous training of learnable parameters.
+Our approach was demonstrated to learn substantially faster than common baselines while achieving competitive predictive performance and showing stable knowledge retention properties.
+Limitations: Depending on hyperparameters like $m$, $\phi$, and $n_{sp}$, PARMESAN can require more computational resources than learned task-heads.
+Moreover, choosing a suitable feature extractor requires taking a trade-off between a general-purpose model versus a data- and task-specific model.
+Open Problems: Possible research opportunities include leveraging self-supervised feature extractors, studying explainability aspects, as well as investigating memory-related topics such as dataset distillation for memory consolidation.
+Inference speed and memory requirements might be further improved, e.g., by using specialized libraries.
+Finally, PARMESAN can be beneficial to other fields such as domain adaptation, transfer learning, and few-shot learning, where flexibility and fast learning are often important.
 Supplementary Material
-Philip Matthias Winter, Maria Wimmer, David Major, Dimitrios Lenis, Astrid Berg, Theresa Neubauer, Gaia Romana De Paolis, Johannes Novotny, Sophia Ulonska, Katja Bühler
+Philip Matthias Winter*, Maria Wimmer*, David Major, Dimitrios Lenis, Astrid Berg, Theresa Neubauer, Gaia Romana De Paolis, Johannes Novotny, Sophia Ulonska, Katja Bühler
 Appendix 0.A Method
@@
 Each node in level $l$ typically has several child nodes in level $l-1$, e.g., $n_{ch}=2^{dim}$.
-Leaf nodes in level $l=0$ refer to memory labels.
+Leaf nodes in level $l=0$ refer to memory labels $z_{j}$.
 We assume that resolutions of $h^{1}$, $x$, and $z$ match, implying that each node in level $l=1$ has exactly one child node in $l=0$.
@@
 We initialize parent similarity scores $s_{k,par}$ with a value of 1.0 for all nodes.
+Some neural architectures may not provide desired feature pyramids in the sense that resolutions are halved at every level or that bottlenecks have small resolutions.
+In these cases, connectivity kernels can either be adjusted accordingly, or missing representations can be inserted by interpolation or extrapolation of neighboring representations.
 Technical Aspects: In order to avoid increasingly exhaustive search due to a rapidly growing number of nodes, we set $\phi$ such that hardware memory requirements remain quasi-constant across levels.
@@
 We implemented all methods and experiments using the PyTorch framework [58].
-We apply the exact same pre- and postprocessing to retain comparability.
+We apply the exact same pre and postprocessing to retain comparability.
 We conduct all experiments on a single NVIDIA GeForce RTX 3090 (24 GB) graphics card.
@@
 For MP, all operations can be done on GPU.
-Appendix 0.B Experiments
+Appendix 0.B Experimental Setup and Additional Experiments
+In this section, we provide more details regarding our overall experimtal setup and perform some additional experiments.
 0.B.1 Datasets
@@
 We rescale all samples to a resolution of 512$\times$1024 pixels for CITY and 512$\times$512 pixels for JSRT.
-For CITY disparity maps, we observe that some pixels near image borders are labeled incorrectly. We manually mask out these erroneous labels and treat them as invalid measurements.
+For CITY disparity maps, we observe that some pixels near image borders are labeled incorrectly.
+We manually mask out these erroneous labels and treat them as invalid measurements.
 Figure 5: Annotations of Cityscapes dataset (original image, depth map, fine semantic map, coarse semantic map).
@@
 0.B.2 Models, Methods, Training and Evaluation
-Models: We use a U-Net [77] variant with residual blocks [30] and instance normalization [96] as our default architecture, resulting in 9 levels (excluding the label-level $l=0$) with respective feature map resolutions of [512[512[ 512, $256$, $128$, $64$, $32$, $16$, $8$, $4$, 2]2]2 ] pixels per dimension.
+Models: We use a U-Net [69] variant with residual blocks [26] and instance normalization [87] as our default architecture, resulting in 9 levels (excluding the label-level $l=0$) with respective feature map resolutions of $[512\times 1024,256\times 512,128\times 256,64\times 128,32\times 64,16\times 32,8\times 16,4\times 8,2\times 4]$ pixels.
 The respective number of feature maps are [6[6[ 6, $12$, $24$, $48$, $96$, $192$, $384$, $768$, 1536]1536]1536 ], resulting in a total of 73.15m learnable parameters.
-For ablations studies, we additionally use a ConvNeXt Tiny with pre-trained weights for feature extraction [50].
-Pre-trained weights were obtained by training ADE20K [112] for semantic segmentation.
-The model comprises 4 levels with feature map resolutions of $[128,64,32,16]$ and corresponding number of feature maps $[96,192,384,786]$.
-To be able to use it with our transduction module, we complete the feature pyramid, i.e., add missing resolutions, via interpolation and extrapolation of neighboring representations.
-We reduce the number of feature maps in high resolutions by a factor of 4 to meet our hardware limitations (see paragraph “Training“ below) and keep the number of feature maps constant for missing lower resolutions.
-This results in $[6,24,96,192,384,768,768,768,768]$ feature maps at the respective resolutions of $[512,256,128,64,32,16,8,4,2]$.
-Method: For memory search, we set $n_{ch}=2^{dim}$ for all levels.
+For some experiments, we use a ConvNeXt Tiny with pre-trained weights for feature extraction [45].
+Pre-training was performed on ADE20K [99] for semantic segmentation.
+We use the publicly available, pre-trained weights provided by [45].
+The ConvNeXt model comprises 4 levels with feature map resolutions of $[128,64,32,16]$ and corresponding number of feature maps $[96,192,384,786]$.
+To be able to use it with our transduction module, we complete feature pyramids, i.e., add missing resolutions, via interpolation and extrapolation of neighboring representations.
+For ConvNeXt, we use the same number of feature maps at respective resolutions as for U-Net.
+Method: For memory search, we set $n_{ch}=4^{dim}$ for all levels.
 For MP, we set $\lambda=1$, $k=16$.
-Convergence is typically achieved within 16 MP steps.
-For TTA, we apply three separate forward passes and down-scale inputs by factors $s=\{0.8,0.9,1\}$, where $s=1$ implies no transformation.
-[…]
+Convergence is typically achieved within 32 MP steps.
+For test-time augmentation (TTA), we use 3 forward passes with down-scaled and reflection-padded variants of $x_{q}$.
+Registration of down-scaled variants with the original variant is achieved by up-scaling and center-cropping the respective predictions.
+We down-scale inputs by factors $s=\{0.8,0.9,1\}$, where $s=1$ implies no transformation.
+A final prediction is obtained by averaging over all passes.
+Training: Some of our experiments (baseline methods and pre-training for PARMESAN) require training of learnable parameters $\theta$.
+We use the cross-entropy loss as the default loss function for semantic segmentation.
+For monocular depth estimation, we regress $d:=\textit{log}\textit{(depth)}$) and use the scale-invariant loss [13] with $\lambda=0.5$.
+We mask out invalid labels both during training and evaluation.
+Training is performed via stochastic gradient descent [66], utilizing the backpropagation algorithm [33, 70] and the AdamW optimizer [34, 47] with default hyperparameters, a minibatch size of 8, a constant learning rate of 0.0001, and a weight decay of 0.00001.
+We apply gradient-norm clipping [57], allowing maximum gradient norms of 50.
+We train our models on randomly cropped patches of 512$\times$512 pixels for CITY and on full images of 512$\times$512 pixels for JSRT.
+For data augmentation during training, we use horizontal flips ($p=0.5$, not for JSRT), vertical flips ($p=0.5$, not for JSRT), rotations ($p=0.5$, ±$\pm 25^{\circ}$), Gaussian noise ($p=0.5$, $\sigma=0.05$), as well as changes in brightness, saturation, and contrast.
 Evaluation: We use exponential moving average models [59, 71, 82] with a rate of 0.99 for evaluation.
-Early stopping [58] with a patience of 3 and a validation verbosity of 3000 update steps is used (400 for JSRT).
+Early stopping [53] with a patience of 5 and a validation verbosity of 3000 update steps is used (400 for JSRT).
 Early stopping in CL experiments only depends on the performance on available classes at the respective CL step.
-We use a validation verbosity of 30 update steps for FSL FT experiments due to small support sets.
+Since classes are highly imbalanced in the 13-1 (7) cl scenario, we apply early stopping using both old and current classes for JOINT and GDUMB.
 We use the Intersection over Union (IoU) [30, 81] as evaluation metric for semantic segmentation, which is defined as the ratio between the size of the intersection and the size of the union of two finite sample sets, measuring their similarity.
@@
 This variant is also referred to as “micro”-averaging.
-Results: We show qualitative prediction results for all ablation studies, CL experiments, and FSL experiments in Fig. 6, Fig. 7, and Fig. 8, respectively.
+Results: We show qualitative prediction results for all ablation studies and CL experiments experiments in Fig.
+6 and Fig.
+7, respectively.
+Experiments A14, A15, and A16 are not mentioned in the main paper text.
+Results in Table 4 indicate moderate memory requirements for our method.
+Performance remains robust in case of low $m$ (A14) or when using less channels (A15, A16).
+Iterative MP and sequentially executed TTA lead to higher $\tau_{i}$ w.r.t.
+our best setup (A1 / A2 / A3).
+Table 4: Size of $M$ (samples stored on GPU), peak allocated GPU memory $\mathit{C}$ at inference, inference speed $\tau_{i}$, and predictive performance on CITY.
+ID Method & Setup $M$ [GB] $\mathit{C}$ [GB] $\tau_{i}$ [s] $\tau_{l}$ [s] $\text{mIoU}_{\text{cl}}$ $\text{mIoU}_{\text{cat}}$
+A0 JOINT - 0.93 0.0068 3.72 82.5 87.4
+A1 ours 6.40 11.0 1.43 0.012 82.1 87.0
+A2 ours, no MP 6.40 10.9 0.25 0.012 81.4 86.4
+A3 ours, no TTA 6.40 10.8 0.48 0.012 81.2 86.3
+A14 ours, $m\!=\!4,\phi\!=\!1$ 0.009 4.44 1.25 0.011 72.8 78.1
+A15 ours, 75% channels 4.82 9.10 1.37 0.011 82.0 86.9
+A16 ours, 50% channels 3.29 7.40 1.32 0.010 81.8 86.7
 0.B.3 Continual Learning Setup
-[…]
-In contrast to image classification, samples can contain classes that belong to different CL steps, which is referred to as overlapping labels.
+The goal in CL is learning a sequence of $N$ tasks $\mathcal{T}=((T_{1},\textit{D}_{1}),...,(T_{N},\textit{D}_{N}))$ on respective datasets $\mathcal{D}=(\textit{D}_{1},...,\textit{D}_{N})$.
+Old data $\textit{D}_{j}$ is inaccessible for CL step $i$ for $j<i$.
+For memory-based CL methods, data from $\textit{D}_{j}$ can still be retained if it is stored in $M$.
+Given a task-solving model (”solver”) $S:X\rightarrow Z$ with inputs $x\in X$, predictions $y\in Z$, and labels $z\in Z$, the objective is minimizing the unbiased sum of losses $\mathbb{E}_{x,z}[L(y,z)]_{\mathcal{T}}$ over all tasks on the entire data distribution.
+We employ two different CL scenarios: 2-2 (4) cat and 13-1 (7) cl.
+At 2-2 (4) cat, data at every CL step comprises labels for all classes of different categories: $\textit{D}_{1}$: void & flat, $\textit{D}_{2}$: construction & object, $\textit{D}_{3}$: nature & sky, and $\textit{D}_{4}$: human & vehicle.
+As proposed by Yang et al.
+[95], the 13-1 (7) cl scenario starts with 13 classes at CL step $i=1$ (i.e., road, sidewalk, building, wall, fence, pole, traffic light, traffic sign, vegetation, terrain, sky, person, rider) and adds one new class at every following CL step (i.e., car, truck, bus, train, motorcycle, bicycle).
+In contrast to image classification, samples for semantic segmentation can contain classes that belong to different CL steps, which is referred to as overlapping labels.
 For the CI-CL scenario, this requires introducing a dedicated background class to avoid data-leakage between different CL steps.
@@
 Although we focus on studying the offline CL scenario in this paper, our method can also be used for online CL, where one does not have access to the entire dataset of a specific task in order to learn continually.
-In our experiments, we apply classical replay (REPLAY) as a baseline method.
-In order to keep labels consistent for replay training, we employ an adaptive masking approach.
+In order to keep labels consistent for REPLAY, we employ an adaptive masking approach.
 For samples of the current CL step $i$, we mask out all background pixels, i.e., pixels which do not belong to currently active classes.
@@
 Training minibatches for REPLAY consist of standard samples in addition to replay samples, which we select randomly from $M$.
-For REPLAY, we observe that using one or more replay samples from $M$ per update step leads to strong memory overfitting, which in turn leads to reduced validation performance.
-We achieve best results with one replay sample every other minibatch.
+For REPLAY, we observe that using 4 or more replay samples from $M$ per update step leads to strong memory overfitting, which in turn leads to reduced validation performance.
+For $m=940$, we achieve best results with minibatches composed of 3 replay samples and 8 standard samples.
+For MiB and REPLAY baselines, the background class is learned in the first CL step. During successive CL steps, background is partially unlearned again depending on new classes.
 Figure 6: Model predictions for all ablation studies with two different query inputs.
 Figure 7: Model predictions for all CL experiments with two different query inputs.
-Figure 8: Model predictions for all FSL experiments with two different query inputs.